### v1

# Equivariant Evidential Deep Learning for Interatomic Potentials

Abstract
Uncertainty quantification (UQ) is critical for assessing the reliability of machine learning interatomic potentials (MLIPs) in molecular dynamics (MD) simulations, identifying extrapolation regimes and enabling uncertainty-aware workflows such as active learning for training dataset construction. Existing UQ approaches for MLIPs are often limited by high computational cost or suboptimal performance. Evidential deep learning (EDL) provides a theoretically grounded single-model alternative that determines both aleatoric and epistemic uncertainty in a single forward pass. However, extending evidential formulations from scalar targets to vector-valued quantities such as atomic forces introduces substantial challenges, particularly in maintaining statistical self-consistency under rotational transformations. To address this, we propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a backbone-agnostic framework that models atomic forces and their uncertainty jointly by representing uncertainty as a full $3\times 3$ symmetric positive definite covariance tensor that transforms equivariantly under rotations. Experiments on diverse molecular benchmarks show that $\text{e}^{2}$IP provides a stronger accuracy-efficiency-reliability balance than the non-equivariant evidential baseline and the widely used ensemble method. It also achieves better data efficiency through the fully equivariant architecture while retaining single-model inference efficiency.
Equivariant Neural Networks, Uncertainty Quantification, Interatomic Potentials, Evidential Deep Learning
1 Introduction
Machine-learning interatomic potentials (MLIPs) enable large-scale molecular dynamics (MD) by approximating ab initio potential energy surfaces at dramatically reduced cost (Behler and Parrinello, 2007). With recent advances in $E(3)$-equivariant neural architectures, MLIPs can now achieve near first-principles accuracy for energies and atomic forces, substantially extending the accessible time- and length-scales of atomistic simulation and supporting a broad range of applications from materials property prediction to non-equilibrium processes (Batzner et al., 2022; Batatia et al., 2022; Thomas et al., 2018; Gasteiger et al., ; Wood et al., 2025; Musaelian et al., 2023). As a result, MLIPs are becoming a key tool for bridging first-principles modeling and large-scale atomistic simulation (Jia et al., 2020; Kozinsky et al., 2023).
Despite these advances, deploying MLIPs reliably in long-timescale MD remains challenging (Perez et al., 2025). Even small force errors can accumulate over millions of integration steps, leading to unphysical trajectories and unreliable conclusions (Fu et al., ). Reliable and computationally efficient uncertainty quantification (UQ) is therefore essential for safe deployment (Bilbrey et al., 2025), as well as for uncertainty-aware workflows such as data selection and adaptive model refinement.
Ensembles are widely used and often regarded as a robust “gold standard” for UQ in MLIPs (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Wollschläger et al., 2023; Willow et al., 2025), but their training and inference costs scale linearly with the number of models, making them expensive for large-scale simulation. Moreover, recent work has raised concerns about the calibration reliability of ensembles in certain settings (Rahaman and others, 2021; Abe et al., 2022; Dern et al., 2024). These limitations motivate single-model UQ approaches that provide uncertainty estimates in a single forward pass. Evidential deep learning (EDL) is a promising direction: by predicting parameters of a prior distribution, it enables direct estimation of both aleatoric and epistemic uncertainty within one pass (Sensoy et al., 2018; Amini et al., 2020; Deng et al., 2023). However, most existing EDL methods target scalar properties (Soleimany et al., 2021). Xu et al. proposed an evidential interatomic potential (eIP) (Xu et al., 2025) for forces, but models uncertainty as independent per-axis variances in a fixed global frame (i.e., a diagonal covariance). While this can express axis-aligned anisotropy, it cannot capture cross-axis correlations (off-diagonal structure) that are often necessary in vector-valued error modeling. More critically, without a full tensor structure, uncertainty cannot transform consistently under rigid rotations, which breaks the coherence of the probabilistic scale and undermines the stability of uncertainty-based ranking decisions across orientations.
Rotation consistency is fundamental for vector-valued force prediction. Forces are equivariant vectors: under a rotation $R$, the target transforms as $\mathbf{F}\mapsto R\mathbf{F}$. When adopting a probabilistic EDL framework for uncertainty modeling, consistency requires not only an equivariant mean prediction but also a distributional structure that transforms accordingly. By basic statistical transformation rules, if a random vector satisfies $\mathbf{Y}^{\prime}=R\mathbf{Y}$, then its covariance must satisfy $\Sigma^{\prime}=R\Sigma R^{\top}$. Therefore, in EDL frameworks where uncertainty is represented by covariance tensors, equivariance of uncertainty is not an additional physical assumption but a requirement for distributional self-consistency.
Motivated by this observation, we propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), an equivariant evidential framework for vector-valued force prediction (Fig. 1). $\text{e}^{2}$IP represents force uncertainty as a full $3\times 3$ symmetric positive definite (SPD) covariance tensor, capturing directional structure and cross-axis correlations while remaining consistent under rotation, and retaining single forward pass efficiency. To guarantee strict positive definiteness and rotation-consistent covariance construction, we parameterize a symmetric matrix in the Lie algebra and map it to the SPD manifold via the matrix exponential. Built on a Normal–Inverse–Wishart (NIW) evidential model, $\text{e}^{2}$IP naturally decomposes uncertainty into aleatoric and epistemic components. To ensure numerical robustness in complex energy landscapes and long-timescale MD, we further introduce equivariance-preserving spectral stabilization and evidence regularization to mitigate overconfidence and prevent numerical degeneracy. Experiments on diverse molecular benchmarks demonstrate that $\text{e}^{2}$IP achieves a better balance of accuracy, efficiency, and reliability compared to baseline methods, providing a practical and physically consistent framework for uncertainty-aware molecular simulations.
Contributions.
  • We introduce $\text{e}^{2}$IP, an $SO(3)$-equivariant evidential framework that models aleatoric and epistemic uncertainty in vector-valued forces as full $3\times 3$ SPD covariances in a single forward pass.
  • We propose an $SO(3)$-consistent full-covariance construction via a Lie-algebra parameterization and matrix exponential, together with equivariance-preserving spectral stabilization for numerically robust training.
  • Through extensive validation across diverse molecular benchmarks, $\text{e}^{2}$IP demonstrates better uncertainty calibration and data efficiency at the single-model inference cost.
Figure 1: e2IP framework. An SE(3)-equivariant GNN predicts the force mean and NIW evidential parameters, and constructs an equivariant SPD covariance $\Sigma_{0}$ (via a symmetric-matrix Lie-algebra head and matrix exponential) to produce tensor-valued uncertainty in a single forward pass.
2 Preliminaries
Problem setup.
We consider an atomic configuration $\mathcal{G}$ with atomic coordinates $\{\vec{x}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ and atomic species $\{z_{i}\in\mathbb{Z}\}_{i=1}^{N}$. Our goal is to predict the atom-wise vector-valued property $\{\vec{y}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$, such as the force acting on atom $i$, while simultaneously estimating the uncertainty associated with each atomic prediction. Within the evidential deep learning framework, the uncertainty for each atom is represented by a $3\times 3$ covariance matrix.
Equivariant backbone features.
We assume an $SE(3)$-equivariant backbone $f_{\theta}$ (e.g., NequIP (Batzner et al., 2022), MACE (Batatia et al., 2022)) that maps $\mathcal{G}$ to node-wise latent features organized in irreducible representation (irrep) spaces. For each atom $i$, the backbone outputs a collection of irrep-valued features $\{\mathbf{h}_{i}^{l}\}$, including scalars ($l=0$) and higher-order tensors ($l>0$). These features will later be used by task-specific heads to produce both the predictive mean and an uncertainty estimate.
2.1 Equivariance requirements for mean and covariance
Let $g=(R,\mathbf{t})\in SE(3)$, where $R$ is an orthogonal matrix with $\det(R)=1$ and $\mathbf{t}\in\mathbb{R}^{3}$. Since atomic forces are translation-invariant, physical consistency requires the predicted target to transform as a vector:
$\vec{\gamma}_{i}(g\cdot\mathcal{G})=R\,\vec{\gamma}_{i}(\mathcal{G}).$ (1)
The uncertainty associated with a vector prediction is represented by a symmetric positive definite (SPD) rank-2 tensor $\mathbf{\Sigma}_{i}\in\mathbb{S}^{3}_{++}$ (e.g., a covariance or scale matrix). Rotation consistency then requires
$\mathbf{\Sigma}_{i}(g\cdot\mathcal{G})=R\,\mathbf{\Sigma}_{i}(\mathcal{G})\,R^{\top}.$ (2)
In particular, predicting independent per-axis variances (a diagonal matrix in a fixed global basis) generally violates Eq. (2) under non-trivial rotations and cannot capture anisotropic correlations aligned with local geometric directions.
2.2 Irrep structure of symmetric rank-2 tensors
Any symmetric matrix $\mathbf{S}_{i}\in\mathbb{S}^{3}$ admits an orthogonal decomposition into an isotropic trace component and a traceless symmetric component:
$\mathbf{S}_{i}=\underbrace{\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}}_{0e}\;+\;\underbrace{\left(\mathbf{S}_{i}-\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}\right)}_{2e}.$ (3)
From the viewpoint of $SO(3)$ representation theory, the space of symmetric rank-2 tensors decomposes as $(0e\oplus 2e)$: the scalar ($0e$) channel corresponds to the trace, and the five-dimensional ($2e$) channel corresponds to the traceless symmetric part. This structure provides a natural, rotation-compatible parameterization for constructing $\mathbf{S}_{i}$ (and hence $\mathbf{\Sigma}_{0,i}$) from equivariant features.
3 Methodology
Overview.
We propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a single forward pass probabilistic framework for atom-wise vector regression with rotation-consistent uncertainty (Fig. 1). Given an input structure, an $SE(3)$-equivariant network jointly predicts the force mean and a full $3\times 3$ SPD uncertainty tensor for each atom. Uncertainty is modeled using a Normal–Inverse–Wishart (NIW) evidential formulation, yielding a multivariate Student-$t$ predictive distribution with explicit aleatoric and epistemic decomposition. The method is built on two key components: (i) an equivariant SPD covariance construction via a symmetric-matrix (Lie algebra) parameterization and matrix exponential, and (ii) an equivariance-preserving spectral stabilizer that bounds covariance scale during training, ensuring strict SPD validity, rotation consistency, and numerical stability for full-covariance learning. For completeness, we summarize all symbols and validity conditions in Appendix B.
3.1 Multivariate Evidential Learning Framework
We model the observation $\vec{y}_{i}$ with a Gaussian likelihood whose mean and covariance are latent:
$p(\vec{y}_{i}\mid\vec{\mu}_{i},\mathbf{\Sigma}_{i})=\mathcal{N}(\vec{y}_{i}\mid\vec{\mu}_{i},\mathbf{\Sigma}_{i}).$ (4)
Following multivariate evidential regression (Meinert and Lavin, 2021), we place a Normal–Inverse–Wishart (NIW) prior over $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$:
$\mathbf{\Sigma}_{i}\sim\mathcal{W}^{-1}(\mathbf{\Psi}_{i},\nu_{i}),\qquad\vec{\mu}_{i}\mid\mathbf{\Sigma}_{i}\sim\mathcal{N}\!\left(\vec{\gamma}_{i},\frac{1}{\kappa_{i}}\mathbf{\Sigma}_{i}\right).$ (5)
Physically interpretable scale parameterization.
To obtain an interpretable scale matrix from the network, we reparameterize
$\mathbf{\Psi}_{i}\triangleq\nu_{i}\mathbf{\Sigma}_{0,i},$ (6)
where $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ is an SPD matrix and describes the shape of the aleatoric noise (up to a scalar factor controlled by $\nu_{i}$).
Posterior predictive.
Marginalizing $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$ yields a multivariate Student-$t$ predictive distribution:
$\begin{split}p(\vec{y}_{i}\mid\vec{\gamma}_{i},\mathbf{\Sigma}_{0,i},\nu_{i},\kappa_{i})=\mathrm{St}_{m_{i}}\!\left(\vec{y}_{i}\,\middle|\,\vec{\gamma}_{i},\;\mathbf{\Lambda}_{i}\right),\\ m_{i}\triangleq\nu_{i}-d+1,\quad\mathbf{\Lambda}_{i}\triangleq\frac{\nu_{i}(\kappa_{i}+1)}{\kappa_{i}(\nu_{i}-d+1)}\mathbf{\Sigma}_{0,i}.\end{split}$ (7)
A full derivation and the expanded negative log-likelihood are given in Appendix C.
Aleatoric and epistemic uncertainty decomposition.
The NIW hierarchy induces an uncertainty decomposition:
$\begin{split}&\mathbf{U}_{\text{ale}}\triangleq\mathbb{E}[\mathbf{\Sigma}_{i}]=\frac{\nu_{i}\mathbf{\Sigma}_{0,i}}{\nu_{i}-d-1}\\ &\mathbf{U}_{\text{epi}}\triangleq\mathrm{Var}[\vec{\mu}_{i}]=\frac{1}{\kappa_{i}}\,\mathbb{E}[\mathbf{\Sigma}_{i}]=\frac{\nu_{i}\mathbf{\Sigma}_{0,i}}{\kappa_{i}(\nu_{i}-d-1)}\end{split}$ (8)
$\mathbf{U}_{\text{ale}}$ reflects observation noise, while $\mathbf{U}_{\text{epi}}$ captures uncertainty in the mean induced by limited data, and both terms transform equivariantly under rotations when $\mathbf{\Sigma}_{0,i}$ does.
Constraints and stable scalar parameterization.
We require $\kappa_{i}>0$ and $\nu_{i}>d+1$ so that $\mathbb{E}[\mathbf{\Sigma}_{i}]$ exists in Eq. (LABEL:eq:ale_epi_main). In practice we enforce a stronger lower bound $\nu_{i}\geq d+2$ to avoid the near-singular regime of $\nu_{i}/(\nu_{i}-d-1)$. We enforce constraints via shifted softplus:
$\begin{split}&\nu_{i}=\mathrm{softplus}(\hat{\nu}_{i})+(d+2)\\ &\kappa_{i}=\mathrm{softplus}(\hat{\kappa}_{i})+\epsilon\end{split}$ (9)
where $\epsilon=10^{-6}$.
3.2 Equivariant SPD covariance via Lie algebra
A central challenge is predicting $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ while preserving rotation equivariance ($\mathbf{\Sigma}_{0,i}\mapsto R\mathbf{\Sigma}_{0,i}R^{\top}$). Standard parameterizations (e.g., Cholesky) can be implementation-dependent and may break equivariance.
Lie-algebra to SPD via matrix exponential.
We predict coefficients in the irreducible representation space $(0e\oplus 2e)$, map them to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathfrak{sym}(3)$, and define
$\mathbf{\Sigma}_{0,i}=\exp(\mathcal{S}_{i}).$ (10)
Because $\mathcal{S}_{i}$ is symmetric, it admits an eigen-decomposition $\mathcal{S}_{i}=U\Lambda U^{\top}$ with real eigenvalues; hence $\exp(\mathcal{S}_{i})=U\exp(\Lambda)U^{\top}$ has strictly positive eigenvalues and is therefore SPD. Moreover, the matrix exponential is conjugation-equivariant: for any $R\in SO(3)$,
$\exp(R\mathcal{S}_{i}R^{\top})=R\,\exp(\mathcal{S}_{i})\,R^{\top},$ (11)
which follows from the power-series definition of $\exp(\cdot)$ and the identity $(R\mathcal{S}_{i}R^{\top})^{k}=R\mathcal{S}_{i}^{k}R^{\top}$. Consequently, $\mathbf{\Sigma}_{0,i}$ transforms equivariantly as a rank-2 tensor. (Shumaylov et al., )
Irrep-to-Cartesian map.
Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the predicted coefficients, where $s_{i}$ is the $l=0$ channel and $\mathbf{t}_{i}$ parameterizes the $l=2$ (symmetric traceless) channel. We use a fixed linear map to obtain $\mathcal{S}_{i}\in\mathfrak{sym}(3)$; details (including the constant map $\mathbf{Q}$ and vectorization convention) are provided in Appendix D.
3.3 Spectral stabilization via coefficient-space damping
While Eq. (10) guarantees geometric validity, large-magnitude $\mathcal{S}_{i}$ may lead to ill-conditioned $\mathbf{\Sigma}_{0,i}$ due to the exponential growth/shrinkage of eigenvalues. To stabilize training without explicit eigendecomposition, we damp the coefficients in representation space before constructing $\mathcal{S}_{i}$.
Equivariance-preserving damping.
For the scalar channel, we apply a smooth two-sided damper $\tilde{s}_{i}=\phi_{\mathrm{low}}(\phi_{\mathrm{high}}(s_{i}))$. For the $l=2$ channel, we damp only the magnitude:
$\tilde{\mathbf{t}}_{i}=\alpha_{i}\,\mathbf{t}_{i},\qquad\alpha_{i}\triangleq\frac{\phi_{\mathrm{high}}(\|\mathbf{t}_{i}\|_{2})}{\|\mathbf{t}_{i}\|_{2}+\epsilon}.$ (12)
Since $\|\mathbf{t}_{i}\|_{2}$ is rotation-invariant, $\alpha_{i}$ is invariant, and the operation preserves equivariance by construction. We then form $\tilde{\mathbf{z}}_{i}=[\tilde{s}_{i},\tilde{\mathbf{t}}_{i}]$, map it to $\mathcal{S}_{i}$, and compute $\mathbf{\Sigma}_{0,i}=\exp(\mathcal{S}_{i})$. The explicit piecewise linear–Tanh form of $\phi_{\mathrm{high}}/\phi_{\mathrm{low}}$ is provided in Appendix E.
3.4 Training objective
The evidential model outputs a nominal covariance $\mathbf{\Sigma}_{0,i}$ and two evidence parameters $(\nu_{i},\kappa_{i})$, optimized by maximum likelihood estimation with a regularization term.
Student-$t$ negative log-likelihood.
Minimizing the negative log-likelihood of Eq. (7) yields the training loss in Eq. (13). For completeness, the expanded form used in the implementation is reported in Appendix C.
$\begin{split}\mathcal{L}_{\mathrm{NLL},i}&=\log\Gamma\!\left(\frac{\nu_{i}-d+1}{2}\right)-\log\Gamma\!\left(\frac{\nu_{i}+1}{2}\right)\\ &\quad+\frac{d}{2}\log\!\left(\frac{\pi\nu_{i}(1+\kappa_{i})}{\kappa_{i}}\right)+\frac{1}{2}\log\left|\mathbf{\Sigma}_{0,i}\right|\\ &\quad+\frac{\nu_{i}+1}{2}\log\!\left(1+\frac{\kappa_{i}}{\nu_{i}(1+\kappa_{i})}\text{M}\right).\\ \\ \text{M}&\triangleq(\vec{y}_{i}-\vec{\gamma}_{i})^{\top}\mathbf{\Sigma}_{0,i}^{-1}(\vec{y}_{i}-\vec{\gamma}_{i})\end{split}$ (13)
Evidence regularization.
To discourage overconfident evidence on hard or out-of-distribution samples, we add a monotonic penalty on evidence:
$\mathcal{L}_{\text{reg,i}}=(\nu_{i}+\kappa_{i})\|\vec{y}_{i}-\vec{\gamma}_{i}\|_{2},$ (14)
and optimize $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{NLL}}+\lambda\mathcal{L}_{\text{reg}}$. Note that $\partial\mathcal{L}_{\text{reg}}/\partial\nu_{i}=\|\vec{y}_{i}-\vec{\gamma}_{i}\|_{2}$, so larger errors induce a stronger gradient opposing evidence accumulation.
Scalar proxy for tensor epistemic uncertainty.
For comparison against scalar error metrics, we report
$u_{\text{scalar}}=\sqrt{\mathrm{tr}(\mathbf{U}_{\text{epi}})/3}.$ (15)
The trace is rotation-invariant ($\mathrm{tr}(R\mathbf{A}R^{\top})=\mathrm{tr}(\mathbf{A})$), providing a consistent scalar summary.
Algorithm 1 Forward and loss on atom $i$
[…]
10: Compute $\mathcal{L}_{\mathrm{NLL},i}$
11: Compute $\mathcal{L}_{\mathrm{reg},i}$
12: Compute $\mathcal{L}_{\mathrm{energy}}$
[…]
14:Return $\mathcal{L}_{i}\leftarrow\lambda_{\mathrm{energy}}\mathcal{L}_{i,\mathrm{energy}}+\lambda_{\mathrm{forces}}\mathcal{L}_{i,\mathrm{forces}}$
4 Experiments
We evaluate $\text{e}^{2}$IP across molecular dynamics benchmarks to characterize both predictive performance and the usability of its uncertainty estimates. Specifically, we study (i) probabilistic reliability and calibration on ab initio liquid water against a deep ensemble baseline, (ii) sample efficiency and geometric consistency on supramolecular MD22 systems with non-equivariant evidential baselines and an equivariance ablation, (iii) generalization under distribution shift and backbone migration on the Out of Distribution (OOD) Silica glass benchmark, and (iv) numerical robustness of full-covariance learning via spectral-stabilization ablations. Unless otherwise stated, we use AlphaNet (Yin et al., 2025) as the $SE(3)$-equivariant backbone.
4.1 Evaluation on Ab Initio Liquid Water Thermodynamics
We evaluate $\text{e}^{2}$IP on the ab initio liquid water benchmark of Cheng et al. (Cheng et al., 2019) (revPBE0-D3). Liquid water is a demanding condensed-phase setting: strong anharmonic fluctuations and nuclear quantum effects induce substantial, configuration-dependent variability in atomic forces. This benchmark stress-tests whether UQ remains reliable and actionable under realistic physical noise, i.e., whether predicted confidence levels agree with empirical frequencies which is a prerequisite for uncertainty-aware workflows such as risk control in long-timescale molecular dynamics, anomaly screening, and ranking decisions in active learning.
We use an 80%/10%/10% train/validation/test split and report mean $\pm$ std over five random seeds. We compare against an ensemble baseline consisting of five independently trained models with different random seeds, reporting both its raw output (Ensemble(raw)) and a post-hoc calibrated variant (Ensemble(cal)). The calibration fits a single isotropic scale $\sigma^{2}$ on the validation set to target $\sim 0.9$ coverage and is fixed for test evaluation (Appendix G).
Figure 2: Uncertainty diagnostics on liquid water. Top-left: reliability diagram (observed vs. nominal coverage; dashed line is ideal). Top-right: PIT CDF deviation $\hat{F}_{U}(p)-p$ (zero is ideal). Ensemble (raw) is strongly overconfident, while post-hoc calibration over-corrects and becomes under-confident; $\text{e}^{2}$IP stays closest to the ideal in both views. Bottom: alignment between force error and predicted epistemic scale, plotting Force RMSE against $\sqrt{\mathrm{tr}(\mathbf{U}_{\text{epi}})/3}$ (both in log10), shown as 2D density (counts in log scale) for $\text{e}^{2}$IP (left) and Ensemble (right). $\text{e}^{2}$IP exhibits a clearer monotonic relationship and higher Spearman correlation (0.78 vs. 0.61), indicating better error-aware ranking.
We report point accuracy (Force MAE), probabilistic quality (Energy Score, ES (Gneiting and Raftery, 2007); negative log-likelihood, NLL), and calibration diagnostics (Coverage@95, Calibration Error, CE, and PIT; Fig. 2) (Table 1). While ES and NLL are both proper scoring rules, NLL is more sensitive to overconfident failures and distributional mis-scaling, whereas ES emphasizes average multivariate distribution matching. PIT assesses distributional calibration: a calibrated model yields uniform PIT values, hence $\hat{F}_{U}(p)-p\approx 0$. The results show a clear accuracy–efficiency–reliability trade-off: the ensemble achieves lower MAE via model averaging but incurs substantially higher inference latency, while $\text{e}^{2}$IP attains competitive accuracy with a single forward pass and yields an approximately $4.5\times$ per-atom speedup, which is crucial for large-scale or long-timescale simulations.
Table 1: Results on the ab initio liquid water benchmark (Cheng et al., 2019). We report Force MAE (lower is better), Energy Score (ES, lower is better), Coverage@95 (closer to $0.95$ is better), Calibration Error (CE, lower is better), Negative log-likelihood (NLL, lower is better), and inference time per atom (ms; lower is better). All numbers are mean $\pm$ std over five random seeds. For the ensemble baseline, the isotropic calibration parameter $\sigma^{2}$ is fit on the validation set and then fixed for test evaluation. Best results are highlighted in bold.
[…]
Ensemble(raw) 0.386 $\pm$ 0.012 0.64 $\pm$ 0.00 0.255 0.380 237.62 0.463 $\pm$ 0.067
Ensemble(cal) 0.386 $\pm$ 0.012 0.67 $\pm$ 0.00 0.931 0.256 2.74 0.463 $\pm$ 0.067
[…]
More importantly, probabilistic and calibration metrics reveal qualitative differences beyond point error. Ensemble(raw) is severely overconfident (reliability/PIT), leading to extremely low Coverage@95 and extremely large NLL, indicating that its raw uncertainty is not directly usable in practice. Ensemble(cal) reduces global overconfidence via isotropic scaling, but a single scalar cannot capture the configuration-dependent and orientation-specific error structure in condensed phases. In contrast, $\text{e}^{2}$IP predicts a full covariance tensor end-to-end, explicitly modeling anisotropic and correlated uncertainty in 3D force space; this yields markedly better probabilistic consistency (lower NLL and CE/PIT) and fewer high-confidence-but-wrong predictions, making it more suitable for uncertainty-weighted simulation. Fig. 2 further shows stronger alignment between predicted epistemic scale and true force error, directly improving ranking-based decisions such as active learning (Kulichenko et al., 2023; Jung et al., 2024).
Although $\text{e}^{2}$IP scores slightly lower on ES and Coverage@95, the high score of Ensemble(cal) depends on calibration, while the uncalibrated Ensemble(raw) only achieves 0.255 on this metric. Additionally, calibration also results in a slight increase in ES for Ensemble. Considering the Spearman correlation, CE, and NLL metrics, where $\text{e}^{2}$IP significantly outperforms ensembles, we find that $\text{e}^{2}$IP offers a more favorable accuracy–efficiency–reliability balance for uncertainty-aware MD workflows.
4.2 Sample Efficiency with Equivariance
We then evaluate sample efficiency on two molecular systems from the MD22 dataset, Buckyball-Catcher ($C_{60}$$C_{60}$) and Double-walled carbon nanotubes (DWCT), which feature long-range and collective interactions and are more challenging than other MD22 molecules. We compare $\text{e}^{2}$IP against two non-equivariant baselines: eIP (Xu et al., 2025) and Ours w/o equiv., an ablation that preserves the same likelihood objective (NLL) and uncertainty regularization but removes rotational equivariance by constructing a non-equivariant $\Sigma_{0}$ via a Cholesky parameterization. All methods are trained using 30%, 50%, 70%, and 100% of the training data. Fig. 3 reports force MAE and Spearman rank correlation as a function of training-set usage, capturing absolute accuracy and ranking quality, respectively.
Across both systems, $\text{e}^{2}$IP consistently exhibits improved data efficiency, with the largest and most uniform gains on DWCT. This behavior aligns with the strong geometric anisotropy and extended interactions in nanotube systems, where enforcing rotational equivariance provides a powerful inductive bias that improves generalization under limited data. Removing equivariance (Ours w/o equiv.) degrades performance despite sharing the same training objective, indicating that the gains are not driven solely by the loss design. On Buckyball-Catcher, $\text{e}^{2}$IP also improves sample efficiency, although the margin is less pronounced. We note that $\text{e}^{2}$IP maintains stable Spearman correlation as data decreases, suggesting that it preserves relative force ordering while improving accuracy. We further probe this inductive bias by evaluating on randomly rotated DWCT test configurations, where equivariance becomes essential and $\text{e}^{2}$IP achieves a larger margin.
Figure 3: Data efficiency on MD22 subsets. Force MAE (top; lower is better) and Spearman rank correlation (bottom; higher is better) on Buckyball-Catcher (left) and DWCT (right) as a function of training-set usage (30/50/70/100%). We compare $\text{e}^{2}$IP with eIP (Xu et al., 2025) and Ours w/o equiv.. $\text{e}^{2}$IP achieves consistently lower MAE across data regimes, with the largest gains on DWCT, while maintaining stable rank correlation.
A physically meaningful force field should be objective under rigid-body rotations: rotating the coordinates by $R$ should rotate the forces by the same $R$. To explicitly probe this property, we evaluate robustness under proper 3D rotations from the special orthogonal group
$SO(3)=\{R\in\mathbb{R}^{3\times 3}\mid R^{\top}R=I,\ \det(R)=+1\}.$
Given a configuration with coordinates of atom $i$ $\vec{x}_{i}$ and reference forces $\vec{f}_{i}$, we construct a rotated input $\vec{x}_{i}^{\prime}=R\vec{x}_{i}$ and the correspondingly rotated labels $\vec{f}_{i}^{\prime}=R\vec{f}_{i}$. We then compute the force error magnitude $e_{i}=\lVert\vec{f}_{i}^{\prime}-\hat{\vec{f}}_{i}(\vec{x}_{i}^{\prime})\rVert$, and an uncertainty score $u_{i}$ obtained from a rotationally invariant scalar proxy of the predicted uncertainty tensor (Eq. (15)). To assess whether uncertainty remains actionable after transformation, we measure the Spearman rank correlation between uncertainty and error, $\rho=\mathrm{Spearman}(\{u_{i}\},\{e_{i}\})$, report correlations on the original test set ($\rho_{\text{orig}}$) and the rotated test set ($\rho_{\text{rot}}$), and summarize the stability via $\Delta\rho=\rho_{\text{rot}}-\rho_{\text{orig}},$ where values closer to zero indicate better preservation of error–uncertainty alignment.
We focus on DWCT, where equivariance is expected to be particularly beneficial due to pronounced geometric anisotropy and extended interactions. For each DWCT test configuration, we sample an independent random $R\in SO(3)$ to form a rotated test set, and compare $\text{e}^{2}$IP against eIP and the non-equivariant ablation Ours w/o equiv.. This setup isolates whether enforcing rotational equivariance improves transformation consistency and, critically, whether it preserves uncertainty-aware ranking under arbitrary 3D rotations.
Table 2 reports $\Delta\rho$ across different training-data regimes. Applying random $SO(3)$ rotations causes a modest degradation in rank alignment for all methods, which is reasonable in practice: even when the target mapping is equivariant, finite numerical precision and common implementation choices (e.g., discretization, neighbor selection, and approximate geometric features) can introduce small rotation-dependent perturbations. We further verify that our method satisfies $SO(3)$ equivariance to numerical precision; see Appendix H for details. Importantly, the non-equivariant baseline eIP exhibits the largest negative shifts, indicating that its uncertainty ranking is most sensitive to rotations. The ablation Ours w/o equiv. is more stable than eIP, suggesting that matching the likelihood objective and uncertainty regularization helps, but it still shows noticeably larger changes than $\text{e}^{2}$IP. Across all data regimes, $\text{e}^{2}$IP consistently achieves the smallest $|\Delta\rho|$ (Table 2), demonstrating that enforcing rotational equivariance is crucial for maintaining uncertainty–error alignment under arbitrary 3D rotations and thus for reliable, uncertainty-driven ranking in downstream use.
Table 2: Change in Spearman correlation under random $SO(3)$ rotations on DWCT. We report $\Delta\rho=\rho_{\text{rot}}-\rho_{\text{orig}}$ (closer to 0 is better). Best results are highlighted in bold.
Method 30% 50% 70% 100% Avg.
eIP (Xu et al., 2025) -0.0856 -0.1454 -0.1302 -0.1418 -0.1258
e2IP (Ours) w/o equiv -0.0754 -0.0384 -0.0449 -0.0432 -0.0505
e2IP (Ours) -0.0470 -0.0235 -0.0244 -0.0008 -0.0239
4.3 Backbone-migration Evaluation on Silica Glass (OOD)
To stress-test the backbone-migration capability of our uncertainty head, we run controlled experiments on the Silica glass benchmark (Tan et al., 2023) using another equivariant backbone PaiNN (Schütt et al., 2021) as the base predictor. We keep the PaiNN backbone and all training hyperparameters fixed across runs, and train different UQ heads with the same random seed to minimize variance from initialization (full settings in Appendix A). A key aspect of this benchmark is that the official test split is intentionally out-of-distribution (OOD): training structures are generated under low temperature and low deformation rate, whereas test structures come from higher-temperature and higher-deformation-rate trajectories. This shift probes whether a UQ method remains informative when generalization is genuinely difficult, rather than under an in-domain split.
Following the dataset protocol, we subsample 1000 structures from the available training pool for training and use the remaining training structures for validation; evaluation is performed on the OOD test set. Table 3 reports force MAE together with the rank correlation between per-structure force RMSE and predicted uncertainty, where higher correlation indicates better uncertainty–error alignment for decision-making. Across both metrics, $\text{e}^{2}$IP improves over the eIP baseline, demonstrating that our uncertainty head transfers effectively to a different backbone and remains useful under a challenging distribution shift.
Table 3: Results on the Silica test set (OOD). We report force MAE and the spearman correlation coefficient between force RMSE and predicted uncertainty (higher is better). Best results are highlighted in bold.
Method Forces MAE $\downarrow$ $\rho$ $\uparrow$
eIP  (Xu et al., 2025) 1.69 0.7603
$\text{e}^{2}$IP (Ours) 1.49 0.7759
4.4 Ablation Study: Necessity of Spectral Stabilization
Figure 4: Spectral stabilization is necessary for numerical robustness. We track the batch-wise mean and min–max band of the eigenvalue ratio $\lambda_{\max}(\Sigma_{0})/\lambda_{\min}(\Sigma_{0})$ during early training. left (w/ damper): the ratio stays bounded and training remains stable. right (w/o damper): the ratio rapidly explodes and becomes erratic; negative values indicate that $\lambda_{\min}(\Sigma_{0})$ turns non-positive due to finite-precision effects, after which Cholesky factorization fails (dashed line).
Our method constructs the base covariance $\Sigma_{0}$ during training, which must remain positive definite and numerically well-conditioned to support stable optimization (e.g., Cholesky factorization). To isolate the role of the proposed damper, we ablate the spectral stabilization while keeping all other components unchanged on the liquid water dataset from Sec. 4.1. The collapse behavior without spectral stabilization is consistently observed across other benchmarks, suggesting that the instability is a general numerical issue.
Figure 4 monitors the batch-wise mean and min–max band of the condition proxy $\lambda_{\max}(\Sigma_{0})/\lambda_{\min}(\Sigma_{0})$ over early training. With damping (left), the ratio may increase during initial adaptation but quickly enters a bounded regime, and the min–max band remains controlled, indicating that the covariance spectrum stays well-conditioned throughout training.
Without damping (right), the spectrum becomes unstable: the eigenvalue ratio grows by orders of magnitude and fluctuates sharply across samples, signaling severe ill-conditioning. Moreover, the ratio can become negative, which implies $\lambda_{\min}(\Sigma_{0})\leq 0$ and therefore a loss of positive definiteness under finite precision. After the first such event (dashed line), Cholesky factorization fails and training diverges. Overall, this ablation shows that spectral stabilization is not merely a convenience but a necessary component for reliable covariance construction in practice.
5 Discussion and Conclusion
Discussion.
We introduced Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a single forward pass evidential framework for uncertainty estimation in vector-valued force fields. By representing predictive uncertainty as a full $3\times 3$ SPD covariance, $\text{e}^{2}$IP captures directional structure and cross-axis correlations that scalarized formulations cannot express. More broadly, our results indicate that geometric structure plays a direct role not only in predictive accuracy, but also in making uncertainty probabilistically interpretable and actionable for downstream decision-making in molecular simulation.
A central design principle of $\text{e}^{2}$IP is to couple evidential learning with $SE(3)$-equivariant rank-2 tensor modeling, ensuring that uncertainty transforms consistently under rotations. This is realized through a Lie-algebraic covariance parameterization, in which a symmetric matrix is mapped to the SPD manifold via the matrix exponential, providing positive definiteness by construction and aligning naturally with equivariant features. Our ablations further highlight an important practical consideration: learning full covariances can lead to numerical instability under finite precision if the spectrum becomes extreme. The proposed spectral damping and evidence regularization mitigate this issue, stabilizing training and improving calibration and error–uncertainty alignment.
Finally, because the uncertainty head operates on equivariant features and outputs structured tensor components, it can be integrated with different equivariant backbones with minimal modification. This backbone-agnostic design lowers the barrier to adopting structured uncertainty modeling in existing MLIP pipelines.
Limitations and Future Work.
Our evaluation primarily targets $SO(3)$ consistency; extending to $O(3)$ (including reflections) and stress-testing under more extreme non-equilibrium conditions are natural next steps. In addition, the Normal-Inverse-Wishart prior may be restrictive under heavy-tailed residuals; richer priors (e.g., mixtures or robust alternatives) could further improve reliability. An important direction is to quantify downstream gains in uncertainty-aware molecular simulation, such as end-to-end active learning loops and stability improvements in long-timescale MD simulations.
Conclusion.
We presented $\text{e}^{2}$IP, an $SE(3)$-equivariant evidential framework that models aleatoric and epistemic uncertainty in atomic forces via full SPD covariances. Across molecular dynamics benchmarks, $\text{e}^{2}$IP consistently outperforms baselines in uncertainty reliability and calibration quality, while achieving competitive predictive accuracy at single-model inference cost.
These results suggest that incorporating symmetry and tensor structure into uncertainty modeling is a principled and practical alternative to ensembling for uncertainty-aware molecular simulation.
Appendix A Additional details for $\text{e}^{2}$IP
Table 4 summarizes the hyperparameters used for the AlphaNet and PAINN backbones in our experiments. Both models share the same optimization strategy and loss weighting scheme.
Table 4: Hyperparameters for AlphaNet and PAINN models used in experiments.
Category Hyperparameter AlphaNet PAINN
Optimization Optimizer AdamW
Learning Rate $2\times 10^{-4}$
Betas $(0.9,0.999)$
Weight Decay $0.0$
Scheduler ReduceLROnPlateau (min)
Factor / Patience $0.85$ / $50$
Architecture Number of Layers 44 66
Cutoff Radius ($r_{c}$) $5.0$ Å $5.0$ Å
Hidden Channels $128$ $128$
Num. Radial Basis $64$ 128
Main $\chi_{1}$ / MP $\chi_{1}$ $32$ / $32$
$\chi_{2}$ 88
Hidden Channels ($\chi$) $96$
Loss & Reg. Energy Weight $1.0$
Force Weight $10000.0$
Evidence Reg. ($\lambda_{\text{reg}}$) $0.1$
Appendix B Notation and validity conditions
Table LABEL:tab:notation_detailed summarizes symbols used in $\text{e}^{2}$IP.
Table 5: Detailed Notation and Physical Interpretations used in $e^{2}$IP.
Symbol Mathematical Name Physical Interpretation & Intuition
Input, Target & Latent Variables
$\vec{y}_{i}\in\mathbb{R}^{3}$ Target Vector The ground truth atomic force derived from DFT calculations on atom $i$.
$\vec{\mu}_{i}\in\mathbb{R}^{3}$ Latent Mean The unknown ”true” mean force. In the evidential framework, this is treated as a random variable following a Normal distribution.
$\Sigma_{i}\in\mathbb{S}_{++}^{3}$ Latent Covariance The unknown ”true” covariance of the aleatoric noise. It is treated as a random variable following an Inverse-Wishart distribution.
$d$ Dimension The dimensionality of the target vector ($d=3$ for atomic forces).
Evidential Parameters (Network Outputs)
$\vec{\gamma}_{i}\in\mathbb{R}^{3}$ Evidential Mean The model’s best prediction for the atomic force. It serves as the mean of the predictive Student-t distribution.
$\Sigma_{0,i}\in\mathbb{S}_{++}^{3}$ Nominal Covariance A symmetric positive definite matrix determining the shape of the aleatoric uncertainty. It captures directional correlations in force noise.
$\nu_{i}\in\mathbb{R}^{+}$ Degrees of Freedom of IW distribution Represents the quantity of evidence supporting the estimated covariance shape $\Sigma_{0,i}$. Higher $\nu_{i}$ means the model is more confident about the aleatoric noise profile.
$\kappa_{i}\in\mathbb{R}^{+}$ Belief Count Represents the quantity of evidence supporting the mean prediction $\vec{\gamma}_{i}$. Higher $\kappa_{i}$ implies lower epistemic uncertainty (i.e., the model has seen ample similar training data).
$\Psi_{i}\in\mathbb{S}_{++}^{3}$ Scale Matrix The scale matrix for the Inverse-Wishart prior, reparameterized as $\Psi_{i}=\nu_{i}\Sigma_{0,i}$ for physical interpretability.
Uncertainty Quantification Measures
$U_{ale}\in\mathbb{S}_{++}^{3}$ Aleatoric Uncertainty Expected data noise ($\mathbb{E}[\Sigma_{i}]$). It captures irreducible variability (e.g., thermal fluctuations) and transforms as a rank-2 tensor.
$U_{epi}\in\mathbb{S}_{++}^{3}$ Epistemic Uncertainty Model uncertainty ($\text{Var}[\vec{\mu}_{i}]$). It captures the lack of knowledge in OOD regions. Computed as $\frac{\nu_{i}\Sigma_{0,i}}{\kappa_{i}(\nu_{i}-d-1)}$.
$m_{i}$ DoF of t-distribution Degrees of freedom for the posterior Student-t distribution, defined as $m_{i}=\nu_{i}-d+1$.
Geometry, Lie Algebra & Optimization
$\mathcal{S}_{i}\in\mathfrak{sym}(3)$ Lie Algebra Element A symmetric matrix in the linear Lie algebra space. The network predicts this unbounded tensor to ensure unconstrained optimization.
$\exp(\mathcal{S}_{i})$ Matrix Exponential The map $\mathfrak{sym}(3)\to\mathbb{S}_{++}^{3}$ that transforms $\mathcal{S}_{i}$ into the valid SPD covariance $\Sigma_{0,i}$, guaranteeing positive definiteness and equivariance by construction.
$\alpha_{i}$ Spectral Damper A scalar coefficient used to damp the magnitude of the traceless part of $\mathcal{S}_{i}$ during training, preventing numerical instability (ill-conditioning).
$\mathcal{L}_{NLL}$ Negative Log-Likelihood The primary loss function based on the multivariate Student-t distribution, encouraging accuracy and calibration.
$\mathcal{L}_{reg}$ Evidence Regularizer A penalty term that discourages the model from assigning high evidence ($\nu_{i},\kappa_{i}$) to data points with large prediction errors.
Validity conditions.
The predictive Student-$t$ has positive degrees of freedom when $m_{i}=\nu_{i}-d+1>0$, i.e., $\nu_{i}>d-1$. In the main text we additionally require $\nu_{i}>d+1$ so that $\mathbb{E}[\mathbf{\Sigma}_{i}]$ exists and the tensor uncertainty decomposition in Eq. (LABEL:eq:ale_epi_main) is well-defined.
Appendix C Derivation of the multivariate Student-t NLL
This appendix derives the expanded negative log-likelihood used in Eq. (13) in the main text, starting from the NIW predictive distribution.
C.1 Predictive Student-t from an NIW prior
Let $\vec{y}_{i}\in\mathbb{R}^{d}$ denote the target vector and let the likelihood be
$p(\vec{y}_{i}\mid\vec{\mu}_{i},\mathbf{\Sigma}_{i})=\mathcal{N}(\vec{y}_{i}\mid\vec{\mu}_{i},\mathbf{\Sigma}_{i}).$ (16)
We place a Normal–Inverse–Wishart (NIW) prior on $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$ with parameters $(\vec{\gamma}_{i},\kappa_{i},\mathbf{\Psi}_{i},\nu_{i})$:
$\mathbf{\Sigma}_{i}\sim\mathcal{W}^{-1}(\mathbf{\Psi}_{i},\nu_{i}),\qquad\vec{\mu}_{i}\mid\mathbf{\Sigma}_{i}\sim\mathcal{N}\!\left(\vec{\gamma}_{i},\frac{1}{\kappa_{i}}\mathbf{\Sigma}_{i}\right),$ (17)
where $\kappa_{i}>0$ and $\nu_{i}>d-1$ so that the predictive degrees of freedom are positive.
Marginalizing $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$ yields a multivariate Student-t predictive distribution:
$p(\vec{y}_{i}\mid\vec{\gamma}_{i},\kappa_{i},\mathbf{\Psi}_{i},\nu_{i})=St_{\nu_{i}-d+1}\!\left(\vec{y}_{i}\,\middle|\,\vec{\gamma}_{i},\,\frac{1+\kappa_{i}}{\kappa_{i}(\nu_{i}-d+1)}\,\mathbf{\Psi}_{i}\right).$ (18)
C.2 NLL in a log-determinant form
An equivalent closed form for the negative log-likelihood is
$\begin{split}\mathcal{L}_{\mathrm{NLL},i}&=\log\Gamma\!\left(\frac{\nu_{i}-d+1}{2}\right)-\log\Gamma\!\left(\frac{\nu_{i}+1}{2}\right)+\frac{d}{2}\log\!\left(\pi\frac{1+\kappa_{i}}{\kappa_{i}}\right)-\frac{\nu_{i}}{2}\log\left|\mathbf{\Psi}_{i}\right|\\ &\quad+\frac{\nu_{i}+1}{2}\log\left|\mathbf{\Psi}_{i}+\frac{\kappa_{i}}{1+\kappa_{i}}(\vec{y}_{i}-\vec{\gamma}_{i})(\vec{y}_{i}-\vec{\gamma}_{i})^{\top}\right|.\end{split}$ (19)
C.3 Reparameterization $\mathbf{\Psi}_{i}=\nu_{i}\mathbf{\Sigma}_{0,i}$
Following the main text (Eq. (6)), define $\mathbf{\Psi}_{i}\triangleq\nu_{i}\mathbf{\Sigma}_{0,i}$ with $\mathbf{\Sigma}_{0,i}$ being a SPD matrix. Substituting into Eq. (19) gives
$\begin{split}\mathcal{L}_{\mathrm{NLL},i}&=\log\Gamma\!\left(\frac{\nu_{i}-d+1}{2}\right)-\log\Gamma\!\left(\frac{\nu_{i}+1}{2}\right)+\frac{d}{2}\log\!\left(\nu_{i}\pi\frac{1+\kappa_{i}}{\kappa_{i}}\right)-\frac{\nu_{i}}{2}\log\left|\mathbf{\Sigma}_{0,i}\right|\\ &\quad+\frac{\nu_{i}+1}{2}\log\left|\mathbf{\Sigma}_{0,i}+\frac{\kappa_{i}}{\nu_{i}(1+\kappa_{i})}(\vec{y}_{i}-\vec{\gamma}_{i})(\vec{y}_{i}-\vec{\gamma}_{i})^{\top}\right|.\end{split}$ (20)
C.4 Determinant lemma and simplification
Define $\vec{v}_{i}\triangleq\vec{y}_{i}-\vec{\gamma}_{i}$ and $\alpha_{i}\triangleq\frac{\kappa_{i}}{\nu_{i}(1+\kappa_{i})}$. We apply the matrix determinant lemma:
$\left|\mathbf{A}+\mathbf{u}\mathbf{u}^{\top}\right|=|\mathbf{A}|\left(1+\mathbf{u}^{\top}\mathbf{A}^{-1}\mathbf{u}\right),\qquad\text{for invertible }\mathbf{A}.$ (21)
Setting $\mathbf{A}=\mathbf{\Sigma}_{0,i}$ and $\mathbf{u}=\sqrt{\alpha_{i}}\,\vec{v}_{i}$ yields
$\left|\mathbf{\Sigma}_{0,i}+\alpha_{i}\vec{v}_{i}\vec{v}_{i}^{\top}\right|=\left|\mathbf{\Sigma}_{0,i}\right|\left(1+\alpha_{i}\,\vec{v}_{i}^{\top}\mathbf{\Sigma}_{0,i}^{-1}\vec{v}_{i}\right).$ (22)
Taking logarithms,
$\log\left|\mathbf{\Sigma}_{0,i}+\alpha_{i}\vec{v}_{i}\vec{v}_{i}^{\top}\right|=\log\left|\mathbf{\Sigma}_{0,i}\right|+\log\!\left(1+\alpha_{i}\,\vec{v}_{i}^{\top}\mathbf{\Sigma}_{0,i}^{-1}\vec{v}_{i}\right).$ (23)
Substituting Eq. (23) into Eq. (20), the $\log|\mathbf{\Sigma}_{0,i}|$ terms combine as
$\frac{\nu_{i}}{2}\log|\mathbf{\Sigma}_{0,i}|+\frac{\nu_{i}+1}{2}\log|\mathbf{\Sigma}_{0,i}|=\frac{1}{2}\log|\mathbf{\Sigma}_{0,i}|.$ (24)
Therefore we obtain the expanded form used in the main text:
$\begin{split}\mathcal{L}_{\mathrm{NLL},i}&=\log\Gamma\!\left(\frac{\nu_{i}-d+1}{2}\right)-\log\Gamma\!\left(\frac{\nu_{i}+1}{2}\right)+\frac{d}{2}\log\!\left(\frac{\pi\nu_{i}(1+\kappa_{i})}{\kappa_{i}}\right)+\frac{1}{2}\log\left|\mathbf{\Sigma}_{0,i}\right|\\ &\quad+\frac{\nu_{i}+1}{2}\log\!\left(1+\frac{\kappa_{i}}{\nu_{i}(1+\kappa_{i})}(\vec{y}_{i}-\vec{\gamma}_{i})^{\top}\mathbf{\Sigma}_{0,i}^{-1}(\vec{y}_{i}-\vec{\gamma}_{i})\right).\end{split}$ (25)
Appendix D Irrep-to-Cartesian map and the constant basis matrix $\mathbf{Q}$
This appendix specifies the implementation-dependent details omitted in Sec. 3.2. Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the coefficients in $(0e\oplus 2e)$. We map $\mathbf{z}_{i}$ to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathfrak{sym}(3)$ using a fixed Clebsch–Gordan basis.
One possible convention.
We first produce a (not-necessarily symmetric) $3\times 3$ matrix $\mathbf{S}_{i}$ by a linear map,
$\mathrm{vec}(\mathbf{S}_{i})=\mathbf{Q}^{\top}\mathbf{z}_{i}^{\top},\qquad\mathbf{Q}\in\mathbb{R}^{9\times 6},$ (26)
where $\mathrm{vec}(\cdot)$ stacks entries in a fixed order (e.g., row-major). We then symmetrize:
$\mathcal{S}_{i}\triangleq\tfrac{1}{2}(\mathbf{S}_{i}+\mathbf{S}_{i}^{\top})\in\mathfrak{sym}(3).$ (27)
(Any equivalent convention that yields $\mathcal{S}_{i}$ transforming as a rank-2 tensor is valid; we report the exact convention used in code for reproducibility.)
Equivariance.
By construction of $\mathbf{Q}$ from CG coefficients of $(1o\otimes 1o)\to(0e\oplus 2e)$, the resulting $\mathcal{S}_{i}$ transforms as $\mathcal{S}_{i}\mapsto R\mathcal{S}_{i}R^{\top}$ for $R\in SO(3)$.
Implementation details.
We utilized the e3nn.io.CartesianTensor method for this construction. Due to performance inefficiencies in the original source code, we re-implemented the transformation method and validated the consistency of our results against the original implementation.
Appendix E Damping functions used for spectral stabilization
For completeness we provide the explicit damping used in Section 3.3.
Linear–Tanh damper.
We define a smooth piecewise function that is linear near the origin and saturates to a ceiling:
$\phi_{\mathrm{high}}(x)=\begin{cases}x,&|x|\leq\tau,\\ \mathrm{sgn}(x)\left[(c-\tau)\tanh\!\left(\frac{|x|-\tau}{c-\tau}\right)+\tau\right],&|x|>\tau,\end{cases}$ (28)
with threshold $\tau$ and ceiling $c$. The lower damper $\phi_{\mathrm{low}}$ is defined analogously to enforce a floor on the isotropic channel. In our expriment, we use $\tau=4$ and $c=5$.
[…]
Figure 5: Linear–Tanh damping function with threshold $\tau=4$ and ceiling $c=5$. The mapping is linear near the origin and smoothly saturates for large inputs.
Coefficient-space application.
Scalar: $\tilde{s}_{i}=\phi_{\mathrm{low}}(\phi_{\mathrm{high}}(s_{i}))$. Vector: $\tilde{\mathbf{t}}_{i}=\alpha_{i}\mathbf{t}_{i}$ with $\alpha_{i}$ defined in Eq. (12).
Appendix F Evaluation Metrics
We evaluate uncertainty quality for vector-valued force prediction using distribution-aware calibration diagnostics together with negative log-likelihood (NLL) and the multivariate Energy Score (ES). Each atom-level force is a 3D vector. For sample $i$, let the target and prediction be $y_{i}\in\mathbb{R}^{3}$ and $\mu_{i}\in\mathbb{R}^{3}$, and define the error
$e_{i}=y_{i}-\mu_{i}\in\mathbb{R}^{3}.$ (29)
Let $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ denote the predicted positive definite matrix used by the model. For the Gaussian baseline, $\Sigma_{i}$ is the predictive covariance. For our Student-$t$ model, we use the scale matrix $\Lambda_{i}$ defined as
$\Lambda_{i}=\frac{(1+\kappa)\nu_{i}}{\kappa(\nu_{i}-d+1)}\Sigma_{i},\qquad m_{i}:=\nu_{i}-d+1.$ (30)
Predictive families.
Our method (e2IP) outputs a multivariate Student-$t$ predictive distribution with effective degrees of freedom $m_{i}$ and scale matrix $\Lambda_{i}$:
$e_{i}\sim St_{m_{i}}\!\left(0,\Lambda_{i}\right).$ (31)
The deep ensemble baseline uses a Gaussian predictive distribution
$e_{i}\sim\mathcal{N}(0,\Sigma_{i}).$ (32)
where $\Sigma_{i}$ is constructed from epistemic covariance plus an isotropic floor as described in the main text.
F.1 Distribution-aware Calibration Curve, Coverage, and Calibration Error
Calibration is assessed via the probability integral transform (PIT). For each method, we map the multivariate residual $e_{i}$ to a scalar $u_{i}\in(0,1)$ such that $u_{i}$ is uniform when the predictive distribution is calibrated.
Gaussian baseline.
For the ensemble Gaussian predictive model, define the squared Mahalanobis distance
$m_{i}^{2}=e_{i}^{\top}\Sigma_{i}^{-1}e_{i}.$ (33)
If $e_{i}\sim\mathcal{N}(0,\Sigma_{i})$ in 3D, then $m_{i}^{2}\sim\chi^{2}_{3}$, and we set
$u_{i}^{\text{Gauss}}=F_{\chi^{2}_{3}}(m_{i}^{2}),$ (34)
where $F_{\chi^{2}_{3}}$ denotes the CDF of $\chi^{2}_{3}$. Under perfect calibration, $u_{i}^{\text{Gauss}}\sim\mathrm{Unif}(0,1)$.
Student-$t$ model (ours).
For our Student-$t$ predictive model in 3D, we again use the quadratic form $m_{i}^{2}=e_{i}^{\top}\Lambda_{i}^{-1}e_{i}$. If $e_{i}\sim t_{\nu_{i}-d+1}(0,\frac{(1+\kappa)\nu}{\kappa(\nu-d+1)}\Sigma_{i})$, then the scaled statistic
$x_{i}=\frac{m_{i}^{2}}{3}$ (35)
follows an $F$ distribution:
$x_{i}\sim F(3,\nu_{i}-d+1).$ (36)
We therefore define the PIT for our method as
$u_{i}^{t}=F_{F(3,\nu_{i}-d+1)}\!\left(\frac{m_{i}^{2}}{3}\right),$ (37)
where $F_{F(3,\nu_{i}-d+1)}(\cdot)$ is the CDF of the $F(3,\nu_{i})$ distribution. Under perfect calibration, $u_{i}^{t}\sim\mathrm{Unif}(0,1)$.
Unified calibration curve.
Given a grid of nominal confidence levels $\mathcal{P}=\{p_{1},\dots,p_{M}\}\subset(0,1)$, we define the empirical (observed) coverage curve for any method as
$\mathrm{Obs}(p)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[u_{i}\leq p],\qquad p\in\mathcal{P},$ (38)
where $u_{i}$ is the PIT scalar computed using the method’s own predictive family (Gaussian or Student-$t$). The ideal calibration curve satisfies $\mathrm{Obs}(p)=p$.
Coverage at a nominal level.
We report $\mathrm{Coverage@}p^{\star}=\mathrm{Obs}(p^{\star})$ for selected nominal levels (e.g., $p^{\star}\in\{0.8,0.9,0.95\}$). Values closer to $p^{\star}$ indicate better calibration at that confidence level.
Calibration error.
We summarize deviation from the ideal curve using an $\ell_{1}$ calibration error over the grid:
$\mathrm{CE}_{\ell_{1}}=\frac{1}{M}\sum_{m=1}^{M}\left|\mathrm{Obs}(p_{m})-p_{m}\right|.$ (39)
F.2 Negative Log-Likelihood (Model-specific)
We report the average negative log-likelihood under each method’s predictive distribution:
$\mathrm{NLL}=-\frac{1}{N}\sum_{i=1}^{N}\log p(e_{i}),$ (40)
where $p(e_{i})$ is the density of $\mathcal{N}(0,\Sigma_{i})$ for the ensemble baseline and the density of $t_{m_{i}}(0,\Lambda_{i})$ for our method. Since these correspond to different distribution families, NLL should be interpreted as a model-specific proper score.
F.3 Energy Score (Multivariate Proper Scoring Rule)
To assess the full multivariate predictive distribution in a distribution-agnostic manner, we compute the Energy Score (ES), a strictly proper scoring rule for multivariate forecasts. For a predictive random variable $X_{i}\sim F_{i}$ and an independent copy $X_{i}^{\prime}$, the Energy Score is
$\mathrm{ES}_{i}=\mathbb{E}\left[\|X_{i}-y_{i}\|_{2}\right]-\frac{1}{2}\mathbb{E}\left[\|X_{i}-X_{i}^{\prime}\|_{2}\right],$ (41)
where $F_{i}$ is the method’s predictive distribution (Gaussian for the ensemble, Student-$t$ for ours). Lower ES indicates better probabilistic predictions. We estimate the expectations via Monte Carlo with $S$ samples:
$\widehat{\mathrm{ES}}_{i}$ $=\frac{1}{S}\sum_{s=1}^{S}\|x_{i}^{(s)}-y_{i}\|_{2}-\frac{1}{2}\cdot\frac{1}{S}\sum_{s=1}^{S}\|x_{i}^{(s)}-\tilde{x}_{i}^{(s)}\|_{2},$ (42)
$x_{i}^{(s)}$ $\overset{\text{i.i.d.}}{\sim}F_{i},\qquad\tilde{x}_{i}^{(s)}\overset{\text{i.i.d.}}{\sim}F_{i}.$ (43)
We report the mean Energy Score over the evaluation set.
Appendix G Ensemble Predictive Covariance and $\sigma^{2}I$ Calibration
For a deterministic ensemble with member predictions $\{\mu_{i}^{(k)}\}_{k=1}^{K}$, we form the ensemble mean
$\mu_{i}\;=\;\frac{1}{K}\sum_{k=1}^{K}\mu_{i}^{(k)},$ (44)
and an epistemic covariance estimate
$\Sigma^{\text{epi}}_{i}\;=\;\frac{1}{K}\sum_{k=1}^{K}\left(\mu_{i}^{(k)}-\mu_{i}\right)\left(\mu_{i}^{(k)}-\mu_{i}\right)^{\top}.$ (45)
To improve calibration, we optionally add an isotropic variance floor
$\Sigma_{i}(\sigma^{2})\;=\;\Sigma^{\text{epi}}_{i}+\sigma^{2}I,$ (46)
where $\sigma^{2}\geq 0$ is fitted on a validation set by matching the observed coverage at a target level $p_{\mathrm{target}}$ (we use $p_{\mathrm{target}}=0.9$). Concretely, we choose $\sigma^{2}$ such that
$\mathrm{Obs}_{\mathrm{val}}^{\sigma^{2}}(p_{\mathrm{target}})\approx p_{\mathrm{target}},$ (47)
and then evaluate all metrics on the test set using the fixed $\sigma^{2}$.
Appendix H Equivariance Validation via Random Rotations
In this section, we provide a numerical validation of the rotational equivariance properties of the proposed model. Specifically, we verify that the predicted forces and uncertainty tensors transform correctly under arbitrary three-dimensional rotations.
H.1 Experimental setup
We randomly select a representative DWCT structure and apply a set of independent random rotations. For each trial, a rotation matrix $R\in\mathrm{SO}(3)$ is sampled uniformly, and the atomic positions are transformed as $\mathbf{x}\mapsto R\mathbf{x}$.
Let $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{N\times 3}$ denote the predicted forces and let $\boldsymbol{U_{epi}}(\mathbf{x})\in\mathbb{R}^{3\times 3}$ denote the predicted uncertainty tensor. For each rotation, we compute the deviations
$\Delta\mathbf{F}=\mathbf{F}(R\mathbf{x})-R\mathbf{F}(\mathbf{x}),$ (48)
and
$\Delta\boldsymbol{U_{epi}}=\boldsymbol{U_{epi}}(R\mathbf{x})-R\boldsymbol{U_{epi}}(\mathbf{x})R^{\top}.$ (49)
The experiment is repeated for 300 randomly sampled rotations. All components of the resulting deviation tensors are collected and flattened to form empirical distributions.
H.2 Results
Figure 6 shows histograms of the force and uncertainty deviations. Both distributions are sharply centered around zero.
For the force predictions, the deviations exhibit a symmetric distribution with small variance, indicating that the vector-valued outputs satisfy rotational equivariance to numerical precision. For the uncertainty predictions, the deviations are even more tightly concentrated around zero, reflecting the stable equivariant behavior of the second-order tensor output.
The observed residual deviations are several orders of magnitude smaller than the typical scale of the model outputs and are attributed to floating-point rounding errors and accumulated numerical precision limits. These results confirm that the proposed model preserves $\mathrm{SO}(3)$ equivariance for both force and uncertainty predictions in practice.
Figure 6: Histograms of deviations under random $\mathrm{SO}(3)$ rotations for a representative DWCT structure. Left: force deviations $\mathbf{F}(R\mathbf{x})-R\mathbf{F}(\mathbf{x})$. Right: uncertainty deviations $\boldsymbol{\Sigma}(R\mathbf{x})-R\boldsymbol{\Sigma}(\mathbf{x})R^{\top}$. Both distributions are sharply peaked at zero, indicating that the model outputs satisfy rotational equivariance up to numerical precision.

### v2

@@
 Abstract
-Uncertainty quantification (UQ) is critical for assessing the reliability of machine learning interatomic potentials (MLIPs) in molecular dynamics (MD) simulations, identifying extrapolation regimes and enabling uncertainty-aware workflows such as active learning for training dataset construction.
-Existing UQ approaches for MLIPs are often limited by high computational cost or suboptimal performance.
-Evidential deep learning (EDL) provides a theoretically grounded single-model alternative that determines both aleatoric and epistemic uncertainty in a single forward pass.
-However, extending evidential formulations from scalar targets to vector-valued quantities such as atomic forces introduces substantial challenges, particularly in maintaining statistical self-consistency under rotational transformations.
-To address this, we propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a backbone-agnostic framework that models atomic forces and their uncertainty jointly by representing uncertainty as a full $3\times 3$ symmetric positive definite covariance tensor that transforms equivariantly under rotations.
-Experiments on diverse molecular benchmarks show that $\text{e}^{2}$IP provides a stronger accuracy-efficiency-reliability balance than the non-equivariant evidential baseline and the widely used ensemble method.
-It also achieves better data efficiency through the fully equivariant architecture while retaining single-model inference efficiency.
+Uncertainty quantification (UQ) is critical for the reliable deployment of machine learning interatomic potentials (MLIPs) in molecular dynamics (MD) simulations.
+Evidential deep learning (EDL) offers a theoretically grounded single-model alternative to expensive ensemble methods, estimating both aleatoric and epistemic uncertainty in a single forward pass.
+However, extending EDL from scalar targets to vector-valued atomic forces introduces a fundamental challenge: the predicted covariance must transform equivariantly under rotations to maintain statistical self-consistency.
+We propose $\text{e}^{2}$IP, a backbone-agnostic evidential framework that represents force uncertainty as a full $3\times 3$ symmetric positive definite (SPD) covariance tensor, constructed via a tangent-space parameterization and Riemannian exponential map that jointly guarantees positive definiteness and $SO(3)$ equivariance.
+Experiments on liquid water, MD22, rMD17, and silica glass benchmarks show that $\text{e}^{2}$IP achieves better calibration (CE, ES, NLL) than deep ensembles — including heteroskedastic variants — while retaining single-model inference efficiency with ${\sim}4.5\times$ speedup.
 Equivariant Neural Networks, Uncertainty Quantification, Interatomic Potentials, Evidential Deep Learning
@@
 Despite these advances, deploying MLIPs reliably in long-timescale MD remains challenging (Perez et al., 2025).
-Even small force errors can accumulate over millions of integration steps, leading to unphysical trajectories and unreliable conclusions (Fu et al., ).
+Force errors can become problematic when simulations enter out-of-distribution configurations, where ML force fields may produce pathological energy or force predictions (Fu et al., ).
 Reliable and computationally efficient uncertainty quantification (UQ) is therefore essential for safe deployment (Bilbrey et al., 2025), as well as for uncertainty-aware workflows such as data selection and adaptive model refinement.
 Ensembles are widely used and often regarded as a robust “gold standard” for UQ in MLIPs (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Wollschläger et al., 2023; Willow et al., 2025), but their training and inference costs scale linearly with the number of models, making them expensive for large-scale simulation.
-Moreover, recent work has raised concerns about the calibration reliability of ensembles in certain settings (Rahaman and others, 2021; Abe et al., 2022; Dern et al., 2024).
-These limitations motivate single-model UQ approaches that provide uncertainty estimates in a single forward pass.
-Evidential deep learning (EDL) is a promising direction: by predicting parameters of a prior distribution, it enables direct estimation of both aleatoric and epistemic uncertainty within one pass (Sensoy et al., 2018; Amini et al., 2020; Deng et al., 2023).
-However, most existing EDL methods target scalar properties (Soleimany et al., 2021).
+Moreover, recent work has suggested that scaling a single model can be equally effective as full ensembling in certain settings, and that ensemble calibration may require careful post-hoc adjustment (Rahaman and others, 2021; Abe et al., 2022; Dern et al., 2024).
+These limitations motivate single-model UQ approaches.
+Evidential deep learning (EDL) is a promising direction: by predicting parameters of a prior distribution, it estimates both aleatoric and epistemic uncertainty in one forward pass (Sensoy et al., 2018; Amini et al., 2020; Deng et al., 2023).
+However, most EDL methods target scalar properties (Soleimany et al., 2021).
 Xu et al.
-proposed an evidential interatomic potential (eIP) (Xu et al., 2025) for forces, but models uncertainty as independent per-axis variances in a fixed global frame (i.e., a diagonal covariance).
-While this can express axis-aligned anisotropy, it cannot capture cross-axis correlations (off-diagonal structure) that are often necessary in vector-valued error modeling.
-More critically, without a full tensor structure, uncertainty cannot transform consistently under rigid rotations, which breaks the coherence of the probabilistic scale and undermines the stability of uncertainty-based ranking decisions across orientations.
-Rotation consistency is fundamental for vector-valued force prediction.
-Forces are equivariant vectors: under a rotation $R$, the target transforms as $\mathbf{F}\mapsto R\mathbf{F}$.
-When adopting a probabilistic EDL framework for uncertainty modeling, consistency requires not only an equivariant mean prediction but also a distributional structure that transforms accordingly.
-By basic statistical transformation rules, if a random vector satisfies $\mathbf{Y}^{\prime}=R\mathbf{Y}$, then its covariance must satisfy $\Sigma^{\prime}=R\Sigma R^{\top}$.
-Therefore, in EDL frameworks where uncertainty is represented by covariance tensors, equivariance of uncertainty is not an additional physical assumption but a requirement for distributional self-consistency.
-Motivated by this observation, we propose Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), an equivariant evidential framework for vector-valued force prediction (Fig.
-1).
-$\text{e}^{2}$IP represents force uncertainty as a full $3\times 3$ symmetric positive definite (SPD) covariance tensor, capturing directional structure and cross-axis correlations while remaining consistent under rotation, and retaining single forward pass efficiency.
-To guarantee strict positive definiteness and rotation-consistent covariance construction, we parameterize a symmetric matrix in the Lie algebra and map it to the SPD manifold via the matrix exponential.
-Built on a Normal–Inverse–Wishart (NIW) evidential model, $\text{e}^{2}$IP naturally decomposes uncertainty into aleatoric and epistemic components.
-To ensure numerical robustness in complex energy landscapes and long-timescale MD, we further introduce equivariance-preserving spectral stabilization and evidence regularization to mitigate overconfidence and prevent numerical degeneracy.
-Experiments on diverse molecular benchmarks demonstrate that $\text{e}^{2}$IP achieves a better balance of accuracy, efficiency, and reliability compared to baseline methods, providing a practical and physically consistent framework for uncertainty-aware molecular simulations.
+proposed eIP (Xu et al., 2025) for forces, but model uncertainty as independent per-axis variances (diagonal covariance in a fixed global frame).
+This cannot capture cross-axis correlations and, more critically, cannot transform consistently under rigid rotations — breaking the coherence of the probabilistic model across orientations.
+This limitation is not merely aesthetic.
+Forces are equivariant vectors ($\mathbf{F}\mapsto R\mathbf{F}$), so by standard transformation rules the covariance must satisfy $\Sigma^{\prime}=R\Sigma R^{\top}$.
+Equivariance of the uncertainty tensor is therefore not an additional physical assumption but a requirement for distributional self-consistency.
+We propose $\text{e}^{2}$IP (Fig.
+1), an equivariant evidential framework that represents force uncertainty as a full $3\times 3$ SPD covariance tensor.
+Positive definiteness and $SO(3)$ equivariance are jointly guaranteed by parameterizing a symmetric matrix in the tangent space of $\mathbb{S}_{++}^{3}$ and mapping it via the Riemannian exponential (Moakher, 2005; Arsigny et al., 2007).
+Built on a Normal–Inverse–Wishart (NIW) evidential model, $\text{e}^{2}$IP decomposes uncertainty into aleatoric and epistemic components in a single forward pass.
+An equivariance-preserving spectral stabilizer prevents the numerical degeneracy that otherwise causes training to diverge.
+Experiments on four molecular benchmarks show that $\text{e}^{2}$IP achieves better calibration and ranking quality than deep ensembles — including heteroskedastic (MVE) and shallow variants — at ${\sim}4.5\times$ lower inference cost.
 Contributions.
   • We introduce $\text{e}^{2}$IP, an $SO(3)$-equivariant evidential framework that models aleatoric and epistemic uncertainty in vector-valued forces as full $3\times 3$ SPD covariances in a single forward pass.
-  • We propose an $SO(3)$-consistent full-covariance construction via a Lie-algebra parameterization and matrix exponential, together with equivariance-preserving spectral stabilization for numerically robust training.
+  • We propose an $SO(3)$-consistent full-covariance construction via a tangent-space parameterization and Riemannian exponential map on $\mathbb{S}_{++}^{3}$, together with equivariance-preserving spectral stabilization for numerically robust training.
   • Through extensive validation across diverse molecular benchmarks, $\text{e}^{2}$IP demonstrates better uncertainty calibration and data efficiency at the single-model inference cost.
 Figure 1: e2IP framework.
-An SE(3)-equivariant GNN predicts the force mean and NIW evidential parameters, and constructs an equivariant SPD covariance $\Sigma_{0}$ (via a symmetric-matrix Lie-algebra head and matrix exponential) to produce tensor-valued uncertainty in a single forward pass.
+An SE(3)-equivariant GNN predicts the force mean and NIW evidential parameters, and constructs an equivariant SPD covariance $\Sigma_{0}$ (via a tangent-space parameterization and Riemannian exponential map) to produce tensor-valued uncertainty in a single forward pass.
 2 Preliminaries
+This section establishes the notation and the two key requirements—equivariance and positive definiteness—that motivate our covariance construction.
 Problem setup.
-We consider an atomic configuration $\mathcal{G}$ with atomic coordinates $\{\vec{x}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ and atomic species $\{z_{i}\in\mathbb{Z}\}_{i=1}^{N}$.
-Our goal is to predict the atom-wise vector-valued property $\{\vec{y}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$, such as the force acting on atom $i$, while simultaneously estimating the uncertainty associated with each atomic prediction.
-Within the evidential deep learning framework, the uncertainty for each atom is represented by a $3\times 3$ covariance matrix.
-Equivariant backbone features.
-We assume an $SE(3)$-equivariant backbone $f_{\theta}$ (e.g., NequIP (Batzner et al., 2022), MACE (Batatia et al., 2022)) that maps $\mathcal{G}$ to node-wise latent features organized in irreducible representation (irrep) spaces.
-For each atom $i$, the backbone outputs a collection of irrep-valued features $\{\mathbf{h}_{i}^{l}\}$, including scalars ($l=0$) and higher-order tensors ($l>0$).
-These features will later be used by task-specific heads to produce both the predictive mean and an uncertainty estimate.
+We consider an atomic configuration $\mathcal{G}$ with coordinates $\{\vec{x}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ and species $\{z_{i}\in\mathbb{Z}\}_{i=1}^{N}$.
+The goal is to predict atom-wise forces $\{\vec{y}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ together with a per-atom $3\times 3$ covariance matrix representing the uncertainty of each force prediction.
+Equivariant backbone.
+We assume an $SE(3)$-equivariant graph neural network $f_{\theta}$ (e.g., NequIP (Batzner et al., 2022), MACE (Batatia et al., 2022)) that maps $\mathcal{G}$ to node-wise features organized by angular momentum order $l$: for each atom $i$, the backbone outputs features $\{\mathbf{h}_{i}^{l}\}$ including scalars ($l=0$) and higher-order tensors ($l>0$), each transforming as an irreducible representation (irrep) of $SO(3)$.
+Task-specific heads then map these features to the force mean and uncertainty parameters.
 2.1 Equivariance requirements for mean and covariance
-Let $g=(R,\mathbf{t})\in SE(3)$, where $R$ is an orthogonal matrix with $\det(R)=1$ and $\mathbf{t}\in\mathbb{R}^{3}$.
-Since atomic forces are translation-invariant, physical consistency requires the predicted target to transform as a vector:
+Let $g=(R,\mathbf{t})\in SE(3)$ with $R\in SO(3)$.
+Forces are equivariant vectors: rotating a molecular configuration rotates the forces accordingly, so the predicted force mean $\vec{\gamma}_{i}\in\mathbb{R}^{3}$ (the model’s best estimate of the force on atom $i$) must satisfy
 $\vec{\gamma}_{i}(g\cdot\mathcal{G})=R\,\vec{\gamma}_{i}(\mathcal{G}).$ (1)
-The uncertainty associated with a vector prediction is represented by a symmetric positive definite (SPD) rank-2 tensor $\mathbf{\Sigma}_{i}\in\mathbb{S}^{3}_{++}$ (e.g., a covariance or scale matrix).
-Rotation consistency then requires
+The same requirement extends to uncertainty.
+If the force uncertainty is represented by a covariance matrix $\mathbf{\Sigma}_{i}\in\mathbb{S}^{3}_{++}$ (the set of $3\times 3$ symmetric positive definite matrices), it must transform as a rank-2 tensor:
 $\mathbf{\Sigma}_{i}(g\cdot\mathcal{G})=R\,\mathbf{\Sigma}_{i}(\mathcal{G})\,R^{\top}.$ (2)
-In particular, predicting independent per-axis variances (a diagonal matrix in a fixed global basis) generally violates Eq.
+Predicting independent per-axis variances (a diagonal matrix in a fixed global basis) violates Eq.
 (2) under non-trivial rotations and cannot capture anisotropic correlations aligned with local geometric directions.
-2.2 Irrep structure of symmetric rank-2 tensors
-Any symmetric matrix $\mathbf{S}_{i}\in\mathbb{S}^{3}$ admits an orthogonal decomposition into an isotropic trace component and a traceless symmetric component:
-$\mathbf{S}_{i}=\underbrace{\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}}_{0e}\;+\;\underbrace{\left(\mathbf{S}_{i}-\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}\right)}_{2e}.$ (3)
-From the viewpoint of $SO(3)$ representation theory, the space of symmetric rank-2 tensors decomposes as $(0e\oplus 2e)$: the scalar ($0e$) channel corresponds to the trace, and the five-dimensional ($2e$) channel corresponds to the traceless symmetric part.
-This structure provides a natural, rotation-compatible parameterization for constructing $\mathbf{S}_{i}$ (and hence $\mathbf{\Sigma}_{0,i}$) from equivariant features.
+This is the central limitation of existing diagonal evidential approaches (Xu et al., 2025).
+2.2 Irrep decomposition of symmetric matrices
+To construct equivariant symmetric matrices from backbone features, we leverage $SO(3)$ representation theory.
+Any $\mathbf{S}_{i}\in\mathrm{sym}(3)$ decomposes into an isotropic (trace) and a traceless part:
+$\mathbf{S}_{i}=\underbrace{\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}}_{\text{scalar }(0e)}\;+\;\underbrace{\left(\mathbf{S}_{i}-\tfrac{1}{3}\mathrm{tr}(\mathbf{S}_{i})\,\mathbf{I}\right)}_{\text{traceless }(2e)}.$ (3)
+The scalar ($0e$) channel captures the isotropic scale (1 degree of freedom), and the five-dimensional traceless ($2e$) channel captures the anisotropic shape.
+This $(0e\oplus 2e)$ structure provides a natural, rotation-compatible parameterization for constructing $\mathbf{S}_{i}$ from equivariant features, which we exploit in Sec.
+3.2.
 3 Methodology
@@
 Given an input structure, an $SE(3)$-equivariant network jointly predicts the force mean and a full $3\times 3$ SPD uncertainty tensor for each atom.
-Uncertainty is modeled using a Normal–Inverse–Wishart (NIW) evidential formulation, yielding a multivariate Student-$t$ predictive distribution with explicit aleatoric and epistemic decomposition.
-The method is built on two key components: (i) an equivariant SPD covariance construction via a symmetric-matrix (Lie algebra) parameterization and matrix exponential, and (ii) an equivariance-preserving spectral stabilizer that bounds covariance scale during training, ensuring strict SPD validity, rotation consistency, and numerical stability for full-covariance learning.
-For completeness, we summarize all symbols and validity conditions in Appendix B.
+Uncertainty is modeled using a Normal–Inverse–Wishart (NIW) evidential formulation (Meinert and Lavin, 2021), yielding a multivariate Student-$t$ predictive distribution with explicit aleatoric and epistemic decomposition.
+The method builds on existing components — SE(3)-equivariant GNN backbones, CG tensor products for coupling irreps, and multivariate evidential regression via NIW — and introduces two novel contributions: (i) to our knowledge, the first full $3\times 3$ equivariant SPD covariance parameterization for evidential force uncertainty, achieved via a tangent-space parameterization and Riemannian exponential map that jointly guarantees positive definiteness, SO(3) equivariance, and stable NIW training; and (ii) an equivariance-preserving spectral stabilizer via coefficient-space damping that bounds eigenvalue magnitude while preserving equivariance, without which training diverges (Fig.
+4).
+Notation.
+We set $d=3$.
+For atom $i$: $\vec{y}_{i}\in\mathbb{R}^{d}$ is the target force vector; $\vec{\gamma}_{i}\in\mathbb{R}^{d}$ is the predicted force mean; $\vec{\mu}_{i}$ and $\mathbf{\Sigma}_{i}$ are latent mean and covariance (random variables under the NIW prior); $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ is the network-predicted nominal covariance (deterministic); $\nu_{i}>0$ and $\kappa_{i}>0$ are evidence parameters controlling covariance and mean uncertainty, respectively.
+$SE(3)$ denotes the special Euclidean group; $\mathbb{S}_{++}^{3}$ the manifold of $3\times 3$ symmetric positive definite matrices; $\mathrm{sym}(3)$ the vector space of $3\times 3$ real symmetric matrices; and “irrep” stands for irreducible representation of $SO(3)$.
+A comprehensive notation table is provided in Appendix B.
 3.1 Multivariate Evidential Learning Framework
-We model the observation $\vec{y}_{i}$ with a Gaussian likelihood whose mean and covariance are latent:
+Having established that force uncertainty requires an equivariant SPD covariance (Sec.
+2.1), we now describe the probabilistic model that jointly represents force predictions and their uncertainty, then show how it decomposes into aleatoric and epistemic components.
+Generative model.
+For each atom $i$, we model the observed force $\vec{y}_{i}$ as drawn from a Gaussian with unknown mean and covariance:
 $p(\vec{y}_{i}\mid\vec{\mu}_{i},\mathbf{\Sigma}_{i})=\mathcal{N}(\vec{y}_{i}\mid\vec{\mu}_{i},\mathbf{\Sigma}_{i}).$ (4)
-Following multivariate evidential regression (Meinert and Lavin, 2021), we place a Normal–Inverse–Wishart (NIW) prior over $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$:
+Rather than estimating $\vec{\mu}_{i}$ and $\mathbf{\Sigma}_{i}$ directly, we treat them as latent random variables and place a conjugate Normal–Inverse–Wishart (NIW) prior (Meinert and Lavin, 2021):
 $\mathbf{\Sigma}_{i}\sim\mathcal{W}^{-1}(\mathbf{\Psi}_{i},\nu_{i}),\qquad\vec{\mu}_{i}\mid\mathbf{\Sigma}_{i}\sim\mathcal{N}\!\left(\vec{\gamma}_{i},\frac{1}{\kappa_{i}}\mathbf{\Sigma}_{i}\right).$ (5)
-Physically interpretable scale parameterization.
-To obtain an interpretable scale matrix from the network, we reparameterize
+Here the network predicts four quantities per atom: the force mean $\vec{\gamma}_{i}$, a nominal covariance $\mathbf{\Sigma}_{0,i}$ (a deterministic matrix that controls the shape of $\mathbf{\Sigma}_{i}$), and two scalar evidence parameters $\nu_{i}$ (covariance evidence) and $\kappa_{i}$ (mean evidence).
+We reparameterize the IW scale matrix as
 $\mathbf{\Psi}_{i}\triangleq\nu_{i}\mathbf{\Sigma}_{0,i},$ (6)
-where $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ is an SPD matrix and describes the shape of the aleatoric noise (up to a scalar factor controlled by $\nu_{i}$).
+so that $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ directly describes the shape of the aleatoric noise.
 Posterior predictive.
-Marginalizing $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$ yields a multivariate Student-$t$ predictive distribution:
+Marginalizing the latent $(\vec{\mu}_{i},\mathbf{\Sigma}_{i})$ yields a closed-form multivariate Student-$t$ predictive distribution (derivation in Appendix C):
 $\begin{split}p(\vec{y}_{i}\mid\vec{\gamma}_{i},\mathbf{\Sigma}_{0,i},\nu_{i},\kappa_{i})=\mathrm{St}_{m_{i}}\!\left(\vec{y}_{i}\,\middle|\,\vec{\gamma}_{i},\;\mathbf{\Lambda}_{i}\right),\\ m_{i}\triangleq\nu_{i}-d+1,\quad\mathbf{\Lambda}_{i}\triangleq\frac{\nu_{i}(\kappa_{i}+1)}{\kappa_{i}(\nu_{i}-d+1)}\mathbf{\Sigma}_{0,i}.\end{split}$ (7)
-A full derivation and the expanded negative log-likelihood are given in Appendix C.
+The degrees of freedom $m_{i}$ control the heaviness of the tails: low $m_{i}$ (little evidence) produces heavy tails reflecting high uncertainty, while $m_{i}\to\infty$ recovers a Gaussian.
 Aleatoric and epistemic uncertainty decomposition.
@@
 $\begin{split}&\mathbf{U}_{\text{ale}}\triangleq\mathbb{E}[\mathbf{\Sigma}_{i}]=\frac{\nu_{i}\mathbf{\Sigma}_{0,i}}{\nu_{i}-d-1}\\ &\mathbf{U}_{\text{epi}}\triangleq\mathrm{Var}[\vec{\mu}_{i}]=\frac{1}{\kappa_{i}}\,\mathbb{E}[\mathbf{\Sigma}_{i}]=\frac{\nu_{i}\mathbf{\Sigma}_{0,i}}{\kappa_{i}(\nu_{i}-d-1)}\end{split}$ (8)
-$\mathbf{U}_{\text{ale}}$ reflects observation noise, while $\mathbf{U}_{\text{epi}}$ captures uncertainty in the mean induced by limited data, and both terms transform equivariantly under rotations when $\mathbf{\Sigma}_{0,i}$ does.
+$\mathbf{U}_{\text{ale}}$ reflects irreducible observation noise (e.g., thermal fluctuations or DFT numerical noise), while $\mathbf{U}_{\text{epi}}$ captures uncertainty in the predicted mean force due to insufficient or unrepresentative training data.
+Both are full $3\times 3$ SPD tensors with physical directional meaning: for example, $\mathbf{U}_{\text{ale}}$ may be anisotropic along a soft vibrational mode.
+Crucially, because $\mathbf{\Sigma}_{0,i}$ is constructed equivariantly (Sec.
+3.2), both $\mathbf{U}_{\text{ale}}$ and $\mathbf{U}_{\text{epi}}$ transform correctly under $SO(3)$ — rotating the molecule rotates the uncertainty ellipsoid accordingly.
 Constraints and stable scalar parameterization.
@@
 where $\epsilon=10^{-6}$.
-3.2 Equivariant SPD covariance via Lie algebra
-A central challenge is predicting $\mathbf{\Sigma}_{0,i}\in\mathbb{S}_{++}^{3}$ while preserving rotation equivariance ($\mathbf{\Sigma}_{0,i}\mapsto R\mathbf{\Sigma}_{0,i}R^{\top}$).
-Standard parameterizations (e.g., Cholesky) can be implementation-dependent and may break equivariance.
-Lie-algebra to SPD via matrix exponential.
-We predict coefficients in the irreducible representation space $(0e\oplus 2e)$, map them to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathfrak{sym}(3)$, and define
+3.2 Equivariant SPD covariance via tangent-space parameterization
+The evidential framework above requires the nominal covariance $\mathbf{\Sigma}_{0,i}$ to be simultaneously (a) symmetric positive definite, (b) equivariant under $SO(3)$, and (c) parameterized in a way amenable to gradient-based optimization.
+Standard parameterizations such as Cholesky decomposition satisfy (a) and (c) but not (b): the triangular factor $L$ in $\mathbf{\Sigma}_{0,i}=LL^{\top}$ does not transform predictably under rotation.
+We address all three requirements with a tangent-space construction.
+From tangent space to SPD via the Riemannian exponential map.
+The manifold $\mathbb{S}_{++}^{3}$ is a Riemannian symmetric space whose tangent space at the identity is $\mathrm{sym}(3)$ (Moakher, 2005; Arsigny et al., 2007).
+The matrix exponential $\exp\colon\mathrm{sym}(3)\to\mathbb{S}_{++}^{3}$ is the Riemannian exponential map — a diffeomorphism that maps any unconstrained symmetric matrix to a valid SPD matrix.
+We predict $(0e\oplus 2e)$ irrep coefficients from the equivariant backbone, map them to a symmetric matrix $\mathcal{S}_{i}\in\mathrm{sym}(3)$ via a fixed Clebsch–Gordan basis (Sec.
+2.2), and define
 $\mathbf{\Sigma}_{0,i}=\exp(\mathcal{S}_{i}).$ (10)
@@
 Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the predicted coefficients, where $s_{i}$ is the $l=0$ channel and $\mathbf{t}_{i}$ parameterizes the $l=2$ (symmetric traceless) channel.
-We use a fixed linear map to obtain $\mathcal{S}_{i}\in\mathfrak{sym}(3)$; details (including the constant map $\mathbf{Q}$ and vectorization convention) are provided in Appendix D.
+We use a fixed linear map to obtain $\mathcal{S}_{i}\in\mathrm{sym}(3)$; details (including the constant map $\mathbf{Q}$ and vectorization convention) are provided in Appendix D.
 3.3 Spectral stabilization via coefficient-space damping
-While Eq.
-(10) guarantees geometric validity, large-magnitude $\mathcal{S}_{i}$ may lead to ill-conditioned $\mathbf{\Sigma}_{0,i}$ due to the exponential growth/shrinkage of eigenvalues.
-To stabilize training without explicit eigendecomposition, we damp the coefficients in representation space before constructing $\mathcal{S}_{i}$.
+The construction above guarantees that $\mathbf{\Sigma}_{0,i}$ is always SPD and equivariant.
+However, in practice the matrix exponential amplifies extreme eigenvalues: if the network predicts large-magnitude $\mathcal{S}_{i}$, the resulting $\mathbf{\Sigma}_{0,i}$ becomes severely ill-conditioned (eigenvalue ratios $10^{5}$), causing numerical failures in the Cholesky factorization required by the NLL loss (Fig.
+4, right).
+To prevent this without breaking equivariance or requiring explicit eigendecomposition, we damp the irrep coefficients before constructing $\mathcal{S}_{i}$.
 Equivariance-preserving damping.
@@
 3.4 Training objective
-The evidential model outputs a nominal covariance $\mathbf{\Sigma}_{0,i}$ and two evidence parameters $(\nu_{i},\kappa_{i})$, optimized by maximum likelihood estimation with a regularization term.
+We now describe how the full pipeline — backbone $\to$ evidential heads $\to$ tangent-space construction $\to$ spectral damping $\to$ matrix exponential — is trained end-to-end.
+Algorithm 1 summarizes the forward pass.
+Given labeled data $\{(\mathcal{G}_{j},\{\vec{y}_{i}^{j}\})\}$ where $\mathcal{G}_{j}$ is an atomic graph and $\vec{y}_{i}^{j}\in\mathbb{R}^{3}$ is the DFT force on atom $i$, the network outputs $(\vec{\gamma}_{i},\mathbf{\Sigma}_{0,i},\nu_{i},\kappa_{i})$ per atom.
+The loss has two components.
 Student-$t$ negative log-likelihood.
-Minimizing the negative log-likelihood of Eq. (7) yields the training loss in Eq. (13). For completeness, the expanded form used in the implementation is reported in Appendix C.
+Minimizing the negative log-likelihood of Eq.
+(7) yields the training loss in Eq.
+(13).
+Intuitively, the $\log|\mathbf{\Sigma}_{0,i}|$ term penalizes overconfident covariance (too small $\mathbf{\Sigma}_{0,i}$ is costly), while the Mahalanobis term M penalizes force prediction errors weighted by the predicted uncertainty.
+For completeness, the expanded form used in the implementation is reported in Appendix C.
 $\begin{split}\mathcal{L}_{\mathrm{NLL},i}&=\log\Gamma\!\left(\frac{\nu_{i}-d+1}{2}\right)-\log\Gamma\!\left(\frac{\nu_{i}+1}{2}\right)\\ &\quad+\frac{d}{2}\log\!\left(\frac{\pi\nu_{i}(1+\kappa_{i})}{\kappa_{i}}\right)+\frac{1}{2}\log\left|\mathbf{\Sigma}_{0,i}\right|\\ &\quad+\frac{\nu_{i}+1}{2}\log\!\left(1+\frac{\kappa_{i}}{\nu_{i}(1+\kappa_{i})}\text{M}\right).\\ \\ \text{M}&\triangleq(\vec{y}_{i}-\vec{\gamma}_{i})^{\top}\mathbf{\Sigma}_{0,i}^{-1}(\vec{y}_{i}-\vec{\gamma}_{i})\end{split}$ (13)
@@
 The trace is rotation-invariant ($\mathrm{tr}(R\mathbf{A}R^{\top})=\mathrm{tr}(\mathbf{A})$), providing a consistent scalar summary.
-Algorithm 1 Forward and loss on atom $i$
+Algorithm 1 Forward pass and loss computation for atom $i$. Steps 1–3 use the shared backbone; steps 4–9 are the evidential uncertainty head; steps 10–14 compute the loss.
 […]
@@
 4 Experiments
-We evaluate $\text{e}^{2}$IP across molecular dynamics benchmarks to characterize both predictive performance and the usability of its uncertainty estimates.
-Specifically, we study (i) probabilistic reliability and calibration on ab initio liquid water against a deep ensemble baseline, (ii) sample efficiency and geometric consistency on supramolecular MD22 systems with non-equivariant evidential baselines and an equivariance ablation, (iii) generalization under distribution shift and backbone migration on the Out of Distribution (OOD) Silica glass benchmark, and (iv) numerical robustness of full-covariance learning via spectral-stabilization ablations.
-Unless otherwise stated, we use AlphaNet (Yin et al., 2025) as the $SE(3)$-equivariant backbone.
+We evaluate $\text{e}^{2}$IP across five experimental settings designed to probe different aspects of uncertainty quality:
+  • Sec. 4.1: Probabilistic reliability on ab initio liquid water, comparing against deep ensembles (MAP and heteroskedastic), shallow ensembles, and eIP.
+  • Sec. 4.2: Sample efficiency and rotation consistency on MD22 supramolecular systems.
+  • Sec. 4.3: Backbone migration and OOD generalization on the Silica glass benchmark.
+  • Sec. 4.4: Standard force-field evaluation on rMD17 molecular benchmarks.
+  • Sec. 4.5: Ablation of spectral stabilization.
+Unless otherwise stated, we use AlphaNet (Yin et al., 2025) as the $SE(3)$-equivariant backbone. Evaluation metrics (ES, CE, NLL, Coverage, Spearman) are defined in Appendix F.
 4.1 Evaluation on Ab Initio Liquid Water Thermodynamics
-We evaluate $\text{e}^{2}$IP on the ab initio liquid water benchmark of Cheng et al.
-(Cheng et al., 2019) (revPBE0-D3).
-Liquid water is a demanding condensed-phase setting: strong anharmonic fluctuations and nuclear quantum effects induce substantial, configuration-dependent variability in atomic forces.
-This benchmark stress-tests whether UQ remains reliable and actionable under realistic physical noise, i.e., whether predicted confidence levels agree with empirical frequencies which is a prerequisite for uncertainty-aware workflows such as risk control in long-timescale molecular dynamics, anomaly screening, and ranking decisions in active learning.
+We evaluate on the ab initio liquid water benchmark (Cheng et al., 2019) (revPBE0-D3), a demanding condensed-phase setting where strong anharmonic fluctuations produce configuration-dependent force variability.
+This stress-tests whether UQ remains reliable under realistic physical noise — a prerequisite for uncertainty-aware workflows such as active learning and risk control.
 We use an 80%/10%/10% train/validation/test split and report mean $\pm$ std over five random seeds.
-We compare against an ensemble baseline consisting of five independently trained models with different random seeds, reporting both its raw output (Ensemble(raw)) and a post-hoc calibrated variant (Ensemble(cal)).
-The calibration fits a single isotropic scale $\sigma^{2}$ on the validation set to target $\sim 0.9$ coverage and is fixed for test evaluation (Appendix G).
+We compare against several baselines: (1) a deep ensemble (DE) consisting of five independently MAP-trained models, reported both raw (DE(raw)) and with post-hoc isotropic calibration (DE(cal), $\sigma^{2}$ fitted on validation set; Appendix G); (2) a heteroskedastic deep ensemble (MVE-DE) where each of K=5 members predicts both force mean and per-component diagonal variance, trained with Gaussian NLL loss following the standard MVE formulation (Scalia et al., 2020; Kellner and Ceriotti, 2024); (3) a shallow ensemble (SE) with shared backbone and K=5 independent output heads (Kellner and Ceriotti, 2024); and (4) the diagonal evidential baseline eIP (Xu et al., 2025).
 Figure 2: Uncertainty diagnostics on liquid water.
@@
 0.61), indicating better error-aware ranking.
-We report point accuracy (Force MAE), probabilistic quality (Energy Score, ES (Gneiting and Raftery, 2007); negative log-likelihood, NLL), and calibration diagnostics (Coverage@95, Calibration Error, CE, and PIT; Fig.
-2) (Table 1).
-While ES and NLL are both proper scoring rules, NLL is more sensitive to overconfident failures and distributional mis-scaling, whereas ES emphasizes average multivariate distribution matching.
-PIT assesses distributional calibration: a calibrated model yields uniform PIT values, hence $\hat{F}_{U}(p)-p\approx 0$.
-The results show a clear accuracy–efficiency–reliability trade-off: the ensemble achieves lower MAE via model averaging but incurs substantially higher inference latency, while $\text{e}^{2}$IP attains competitive accuracy with a single forward pass and yields an approximately $4.5\times$ per-atom speedup, which is crucial for large-scale or long-timescale simulations.
+Table 1 reports point accuracy (Force MAE), probabilistic quality (Energy Score ES (Gneiting and Raftery, 2007), NLL), and calibration diagnostics (Coverage@95, Calibration Error CE, PIT; Fig.
+2).
+The results show a clear accuracy–efficiency–reliability trade-off: DE achieves the lowest MAE via model averaging but incurs ${\sim}4.5\times$ higher inference cost, while $\text{e}^{2}$IP attains competitive accuracy in a single forward pass.
 Table 1: Results on the ab initio liquid water benchmark (Cheng et al., 2019).
@@
 […]
-Ensemble(raw) 0.386 $\pm$ 0.012 0.64 $\pm$ 0.00 0.255 0.380 237.62 0.463 $\pm$ 0.067
-Ensemble(cal) 0.386 $\pm$ 0.012 0.67 $\pm$ 0.00 0.931 0.256 2.74 0.463 $\pm$ 0.067
-[…]
+DE(raw) 0.386 $\pm$ 0.012 0.64 $\pm$ 0.00 0.255 0.380 237.62 0.463 $\pm$ 0.067
+DE(cal) 0.386 $\pm$ 0.012 0.67 $\pm$ 0.00 0.931 0.256 2.74 0.463 $\pm$ 0.067
+MVE-DE 0.492 0.73 0.970 0.139 1.59 0.463 $\pm$ 0.067
+SE(cal) 0.517 $\pm$ 0.012 0.93 $\pm$ 0.06 0.790 $\pm$ 0.071 0.134 $\pm$ 0.028 15.2 $\pm$ 5.6 0.234 $\pm$ 0.004
+eIP (Xu et al., 2025) 0.499 $\pm$ 0.016 2.15 $\pm$ 2.38 0.652 $\pm$ 0.008 0.314 $\pm$ 0.008 3.13 $\pm$ 0.19 0.094 $\pm$ 0.003
+$\text{e}^{2}$IP (Ours) 0.471 $\pm$ 0.010 0.68 $\pm$ 0.02 0.910 $\pm$ 0.024 0.114 $\pm$ 0.028 1.31 $\pm$ 0.03 0.102 $\pm$ 0.016
 More importantly, probabilistic and calibration metrics reveal qualitative differences beyond point error.
-Ensemble(raw) is severely overconfident (reliability/PIT), leading to extremely low Coverage@95 and extremely large NLL, indicating that its raw uncertainty is not directly usable in practice.
-Ensemble(cal) reduces global overconfidence via isotropic scaling, but a single scalar cannot capture the configuration-dependent and orientation-specific error structure in condensed phases.
+DE(raw) is severely overconfident (reliability/PIT), leading to extremely low Coverage@95 and extremely large NLL, indicating that its raw uncertainty is not directly usable in practice.
+DE(cal) reduces global overconfidence via isotropic scaling, but a single scalar cannot capture the configuration-dependent and orientation-specific error structure in condensed phases.
 In contrast, $\text{e}^{2}$IP predicts a full covariance tensor end-to-end, explicitly modeling anisotropic and correlated uncertainty in 3D force space; this yields markedly better probabilistic consistency (lower NLL and CE/PIT) and fewer high-confidence-but-wrong predictions, making it more suitable for uncertainty-weighted simulation.
@@
 2 further shows stronger alignment between predicted epistemic scale and true force error, directly improving ranking-based decisions such as active learning (Kulichenko et al., 2023; Jung et al., 2024).
-Although $\text{e}^{2}$IP scores slightly lower on ES and Coverage@95, the high score of Ensemble(cal) depends on calibration, while the uncalibrated Ensemble(raw) only achieves 0.255 on this metric.
-Additionally, calibration also results in a slight increase in ES for Ensemble.
-Considering the Spearman correlation, CE, and NLL metrics, where $\text{e}^{2}$IP significantly outperforms ensembles, we find that $\text{e}^{2}$IP offers a more favorable accuracy–efficiency–reliability balance for uncertainty-aware MD workflows.
+To address whether $\text{e}^{2}$IP’s advantage stems from modeling input-dependent aleatoric uncertainty, we additionally compare against MVE-DE, a heteroskedastic deep ensemble where each member outputs both force mean and diagonal variance.
+Despite having the capacity to model aleatoric uncertainty, MVE-DE achieves worse MAE (0.492) than $\text{e}^{2}$IP (0.471) and MAP-DE (0.386), and does not outperform $\text{e}^{2}$IP on any UQ metric (ES 0.73 vs 0.68; CE 0.139 vs 0.114; NLL 1.59 vs 1.31).
+The shallow ensemble (SE) performs even worse (NLL 15.2), suggesting that output-layer diversification alone is insufficient for reliable multivariate force UQ.
+We note that NLL uses different distributional forms: Gaussian for DE/MVE-DE/SE, and Student-$t$ for $\text{e}^{2}$IP (the native predictive from NIW marginalization).
+The Energy Score (ES) is a strictly proper scoring rule that depends only on samples and requires no parametric assumptions (Gneiting and Raftery, 2007), making it directly comparable across distributional families.
+$\text{e}^{2}$IP achieves the best ES across all methods.
+Overall, combining stronger probabilistic reliability (NLL, CE, ES), better error–uncertainty ranking alignment, and substantially lower inference cost (${\sim}4.5\times$ speedup over DE), $\text{e}^{2}$IP provides more actionable uncertainty while retaining competitive accuracy on ab initio liquid water.
 4.2 Sample Efficiency with Equivariance
@@
 On Buckyball-Catcher, $\text{e}^{2}$IP also improves sample efficiency, although the margin is less pronounced.
-We note that $\text{e}^{2}$IP maintains stable Spearman correlation as data decreases, suggesting that it preserves relative force ordering while improving accuracy.
-We further probe this inductive bias by evaluating on randomly rotated DWCT test configurations, where equivariance becomes essential and $\text{e}^{2}$IP achieves a larger margin.
+We note that the non-equivariant variant can match or exceed $\text{e}^{2}$IP on raw accuracy at high data fractions on some systems; understanding when and why this occurs is an interesting open question.
+Nevertheless, $\text{e}^{2}$IP maintains stable Spearman correlation as data decreases, suggesting that it preserves relative force ordering while improving accuracy.
+Importantly, the primary advantage of equivariance is not raw MAE dominance at every data point, but rotation-consistent uncertainty: we further probe this by evaluating on randomly rotated DWCT test configurations (Table 2), where $\text{e}^{2}$IP achieves substantially smaller $|\Delta\rho|$.
 Figure 3: Data efficiency on MD22 subsets.
@@
 Table 2: Change in Spearman correlation under random $SO(3)$ rotations on DWCT.
-We report $\Delta\rho=\rho_{\text{rot}}-\rho_{\text{orig}}$ (closer to 0 is better).
+We report $|\Delta\rho|=|\rho_{\text{rot}}-\rho_{\text{orig}}|$ (closer to 0 is better).
+For multi-seed methods we report mean $\pm$ std.
 Best results are highlighted in bold.
 Method 30% 50% 70% 100% Avg.
-eIP (Xu et al., 2025) -0.0856 -0.1454 -0.1302 -0.1418 -0.1258
-e2IP (Ours) w/o equiv -0.0754 -0.0384 -0.0449 -0.0432 -0.0505
-e2IP (Ours) -0.0470 -0.0235 -0.0244 -0.0008 -0.0239
+DE (K=5) 0.140
+eIP (Xu et al., 2025) 0.0856 0.1454 0.1302 $7±\pm0.0$44 0.1258
+e2IP (Ours) w/o equiv 0.0754 0.0384 0.0449 0.0432 0.0505
+[…]
 4.3 Backbone-migration Evaluation on Silica Glass (OOD)
@@
 Following the dataset protocol, we subsample 1000 structures from the available training pool for training and use the remaining training structures for validation; evaluation is performed on the OOD test set.
-Table 3 reports force MAE together with the rank correlation between per-structure force RMSE and predicted uncertainty, where higher correlation indicates better uncertainty–error alignment for decision-making.
-Across both metrics, $\text{e}^{2}$IP improves over the eIP baseline, demonstrating that our uncertainty head transfers effectively to a different backbone and remains useful under a challenging distribution shift.
-Table 3: Results on the Silica test set (OOD).
-We report force MAE and the spearman correlation coefficient between force RMSE and predicted uncertainty (higher is better).
-Best results are highlighted in bold.
-Method Forces MAE $\downarrow$ $\rho$ $\uparrow$
-eIP  (Xu et al., 2025) 1.69 0.7603
-$\text{e}^{2}$IP (Ours) 1.49 0.7759
+Table 3 reports force MAE, rank correlation, and inference time.
+We additionally trained a DE baseline (K=5 PaiNN models) and re-ran eIP (5 seeds) and $\text{e}^{2}$IP (4 seeds) with error bars.
+DE achieves the lowest MAE at $4.6\times$ inference cost (0.242 vs 0.053 ms/atom).
+$\text{e}^{2}$IP outperforms eIP on both MAE and Spearman, and achieves comparable Spearman to DE (0.775 vs 0.790) while maintaining single-model efficiency.
+This demonstrates that our uncertainty head transfers effectively to a different backbone and remains useful under a challenging distribution shift.
+Table 3: Results on the Silica test set (OOD). Force MAE, Spearman $\rho$, and inference time. Mean $\pm$ std over seeds. Best in bold.
+[…]
+[…]
+[…]
+[…]
+4.4 Evaluation on rMD17 Molecular Force Fields
+To evaluate on standard molecular force-field benchmarks with diverse molecular sizes, we compare DE, eIP, $\text{e}^{2}$IP, and MVE-DE on two molecules from the revised MD17 dataset (Christensen and von Lilienfeld, 2020): aspirin (21 atoms) and azobenzene (24 atoms).
+All methods use the same AlphaNet backbone and hyperparameters.
+Results are reported over multiple seeds (Table 4).
+Table 4: Results on rMD17. MAE, ES, CE, NLL. Mean $\pm$ std. Best in bold.
+[…]
+Asp. DE(cal) .187 .703 .423 3.30
+MVE-DE .496 .830 .444 3.35
+[…]
+[…]
+Azo. DE(cal) .139 .341 .327 2.34
+MVE-DE .640 1.238 .472 4.93
+[…]
+[…]
+On both molecules, $\text{e}^{2}$IP achieves the best CE, ES, and NLL.
+On azobenzene, $\text{e}^{2}$IP also achieves the lowest MAE (0.128 vs DE’s 0.139), demonstrating that the evidential framework can compete with ensembles on point accuracy.
+MVE-DE performs substantially worse than both MAP-DE and $\text{e}^{2}$IP across all metrics, with significantly degraded MAE (0.496 on aspirin, 0.640 on azobenzene) and poor calibration (CE >> 0.44).
+Furthermore, some MVE-DE seeds failed to converge during training (aspirin seed 1, azobenzene seeds 3–4), indicating training instability when applying heteroskedastic Gaussian NLL to autograd-derived forces — a known challenge in MLIP UQ.
+These results, combined with the liquid water experiments, consistently demonstrate that $\text{e}^{2}$IP’s NIW evidential framework provides a more effective joint force–uncertainty parameterization than heteroskedastic ensembles.
 4.5 Ablation Study: Necessity of Spectral Stabilization
@@
 Discussion.
-We introduced Equivariant Evidential Deep Learning for Interatomic Potentials ($\text{e}^{2}$IP), a single forward pass evidential framework for uncertainty estimation in vector-valued force fields.
-By representing predictive uncertainty as a full $3\times 3$ SPD covariance, $\text{e}^{2}$IP captures directional structure and cross-axis correlations that scalarized formulations cannot express.
-More broadly, our results indicate that geometric structure plays a direct role not only in predictive accuracy, but also in making uncertainty probabilistically interpretable and actionable for downstream decision-making in molecular simulation.
-A central design principle of $\text{e}^{2}$IP is to couple evidential learning with $SE(3)$-equivariant rank-2 tensor modeling, ensuring that uncertainty transforms consistently under rotations.
-This is realized through a Lie-algebraic covariance parameterization, in which a symmetric matrix is mapped to the SPD manifold via the matrix exponential, providing positive definiteness by construction and aligning naturally with equivariant features.
-Our ablations further highlight an important practical consideration: learning full covariances can lead to numerical instability under finite precision if the spectrum becomes extreme.
-The proposed spectral damping and evidence regularization mitigate this issue, stabilizing training and improving calibration and error–uncertainty alignment.
-Finally, because the uncertainty head operates on equivariant features and outputs structured tensor components, it can be integrated with different equivariant backbones with minimal modification.
-This backbone-agnostic design lowers the barrier to adopting structured uncertainty modeling in existing MLIP pipelines.
+$\text{e}^{2}$IP couples evidential learning with $SE(3)$-equivariant rank-2 tensor modeling: a tangent-space parameterization maps unconstrained symmetric matrices to SPD covariances via the Riemannian exponential, guaranteeing positive definiteness and rotation consistency by construction.
+The result is a full $3\times 3$ uncertainty tensor that captures directional structure inaccessible to diagonal formulations.
+Our ablations show that spectral damping is essential for stable full-covariance learning (Fig.
+4), and our comparison with heteroskedastic deep ensembles (MVE-DE) demonstrates that the NIW framework achieves superior calibration without the training instability observed in Gaussian NLL-based approaches.
+Because the uncertainty head operates on equivariant features and outputs structured tensor components, it integrates with different equivariant backbones (AlphaNet, PaiNN) with minimal modification.
 Limitations and Future Work.
@@
 $m_{i}$ DoF of t-distribution Degrees of freedom for the posterior Student-t distribution, defined as $m_{i}=\nu_{i}-d+1$.
-Geometry, Lie Algebra & Optimization
-$\mathcal{S}_{i}\in\mathfrak{sym}(3)$ Lie Algebra Element A symmetric matrix in the linear Lie algebra space. The network predicts this unbounded tensor to ensure unconstrained optimization.
-$\exp(\mathcal{S}_{i})$ Matrix Exponential The map $\mathfrak{sym}(3)\to\mathbb{S}_{++}^{3}$ that transforms $\mathcal{S}_{i}$ into the valid SPD covariance $\Sigma_{0,i}$, guaranteeing positive definiteness and equivariance by construction.
+Geometry, Tangent Space & Optimization
+$\mathcal{S}_{i}\in\mathrm{sym}(3)$ Tangent Vector A symmetric matrix in the tangent space of $\mathbb{S}_{++}^{3}$ at the identity.
+The network predicts this unbounded tensor to ensure unconstrained optimization.
+$\exp(\mathcal{S}_{i})$ Riemannian Exponential Map The map $\mathrm{sym}(3)\to\mathbb{S}_{++}^{3}$ that transforms $\mathcal{S}_{i}$ into the valid SPD covariance $\Sigma_{0,i}$, guaranteeing positive definiteness and equivariance by construction (Moakher, 2005; Arsigny et al., 2007).
 $\alpha_{i}$ Spectral Damper A scalar coefficient used to damp the magnitude of the traceless part of $\mathcal{S}_{i}$ during training, preventing numerical instability (ill-conditioning).
@@
 Let $\mathbf{z}_{i}=[s_{i},\mathbf{t}_{i}]\in\mathbb{R}^{1+5}$ denote the coefficients in $(0e\oplus 2e)$.
-We map $\mathbf{z}_{i}$ to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathfrak{sym}(3)$ using a fixed Clebsch–Gordan basis.
+We map $\mathbf{z}_{i}$ to a Cartesian symmetric matrix $\mathcal{S}_{i}\in\mathrm{sym}(3)$ using a fixed Clebsch–Gordan basis.
 One possible convention.
 We first produce a (not-necessarily symmetric) $3\times 3$ matrix $\mathbf{S}_{i}$ by a linear map,
-$\mathrm{vec}(\mathbf{S}_{i})=\mathbf{Q}^{\top}\mathbf{z}_{i}^{\top},\qquad\mathbf{Q}\in\mathbb{R}^{9\times 6},$ (26)
+$\mathrm{vec}(\mathbf{S}_{i})=\mathbf{Q}^{\top}\mathbf{z}_{i},\qquad\mathbf{Q}\in\mathbb{R}^{9\times 6},$ (26)
 where $\mathrm{vec}(\cdot)$ stacks entries in a fixed order (e.g., row-major). We then symmetrize:
-$\mathcal{S}_{i}\triangleq\tfrac{1}{2}(\mathbf{S}_{i}+\mathbf{S}_{i}^{\top})\in\mathfrak{sym}(3).$ (27)
+$\mathcal{S}_{i}\triangleq\tfrac{1}{2}(\mathbf{S}_{i}+\mathbf{S}_{i}^{\top})\in\mathrm{sym}(3).$ (27)
 (Any equivalent convention that yields $\mathcal{S}_{i}$ transforming as a rank-2 tensor is valid; we report the exact convention used in code for reproducibility.)
@@
 $\phi_{\mathrm{high}}(x)=\begin{cases}x,&|x|\leq\tau,\\ \mathrm{sgn}(x)\left[(c-\tau)\tanh\!\left(\frac{|x|-\tau}{c-\tau}\right)+\tau\right],&|x|>\tau,\end{cases}$ (28)
-with threshold $\tau$ and ceiling $c$. The lower damper $\phi_{\mathrm{low}}$ is defined analogously to enforce a floor on the isotropic channel. In our expriment, we use $\tau=4$ and $c=5$.
+with threshold $\tau$ and ceiling $c$. The lower damper $\phi_{\mathrm{low}}$ is defined analogously to enforce a floor on the isotropic channel. In our experiments, we use $\tau=4$ and $c=5$.
 […]